### v1

# Model-agnostic Coreset Selection via LLM-based Concept Bottlenecks

Abstract
Coreset Selection (CS) identifies a subset of training data that achieves model performance comparable to using the entire dataset. Many state-of-the-art CS methods, select coresets using scores whose computation requires training the downstream model on the entire dataset and recording changes in its behavior on samples as it trains (training dynamics). These scores are inefficient to compute and hard to interpret as they do not indicate whether a sample is difficult to learn in general or only for a specific model. Our work addresses these challenges by proposing an interpretable score that gauges a sample’s difficulty using human-understandable textual attributes (concepts) independent of any downstream model. Specifically, we measure the alignment between a sample’s visual features and concept bottlenecks, derived via large language models, by training a linear concept bottleneck layer and compute the sample’s difficulty score using it. We then use this score and a stratified sampling strategy to identify the coreset. Crucially, our score is efficiently computable without training the downstream model on the full dataset even once, leads to high-performing coresets for various downstream models, and is computable even for an unlabeled dataset. Through experiments on CIFAR-10, CIFAR-100, and ImageNet-1K, we show our coresets outperform random subsets, even at high pruning rates, and achieve model performance comparable to or better than coresets found by training dynamics-based methods.
Machine Learning, ICML
1 Introduction
Machine learning (ML) pipelines are becoming increasingly intensive regarding their data and compute requirements (Touvron et al., 2023; Achiam et al., 2023). While in line with the empirical neural scaling laws (Kaplan et al., 2020; Hestness et al., 2017; Henighan et al., 2020; Rosenfeld et al., 2019) where a model’s performance improves with the increasing model and training data size, these improvements come at an unsustainable cost of compute/energy. Recently, (Sorscher et al., 2022) showed analytically that data pruning can enable models to achieve an exponential reduction in the test error with an increasing dataset size.
Coreset Selection (CS) (Mirzasoleiman et al., 2020; Guo et al., 2022; Paul et al., 2021; Xia et al., 2022; Maharana et al., 2023; Zheng et al., 2022; Choi et al., 2024) aims to improve the efficiency of training large ML models by pruning a large dataset and retaining only a small subset of representative samples. Most CS methods, use metrics to estimate the difficulty/importance of every training sample and then use a sampling strategy to produce a coreset, that leads to similar performance on a downstream model as training it on the entire dataset. Many state-of-the-art (SOTA) CS methods use the downstream model’s training dynamics — the model’s behavior on a sample during training, to generate a difficulty score for each sample. This enables accurate estimation of the data difficulty but requires training the downstream model on the entire dataset at least once, which can be costly when training a large model on a large dataset. While (Coleman et al., 2019) showed that a coreset selected using training dynamics of a mid-sized proxy model (eg., ResNet-18) is effective for a larger downstream model (eg., ResNet-50), training even such a model is not feasible for large datasets. Moreover, these scores are difficult to interpret as they are downstream model-dependent and cannot inform if the sample will be hard or easy to learn for a different downstream model (without training it first). Thus, we aim to answer the following question “How to efficiently estimate a sample’s difficulty in an interpretable and model-agnostic way i.e. without the knowledge of the architecture or training dynamics of the downstream model?”
Figure 1: : Overview of our approach: We start by prompting the language models to generate concept annotation for $N$ class label names in the dataset and select $k$ most discriminative attributes (per class) to form the concept bottleneck, which are passed through a text encoder ($\mathcal{T}_{enc}$) to obtain the bottleneck embedding matrix ($E_{C}$). The visual information of a training sample $x$ extracted via the visual encoder ($\mathcal{V}_{enc}$) is then aligned to the embedding matrix using a linear concept bottleneck layer trained for $T$ epochs. Our difficulty score for $x$ is then computed as the average margin (i.e., the difference between the likelihood of the correct and other classes) over $T$ epochs. Finally, a coreset is selected using stratified sampling which is then used to train an unknown downstream model.
To address this question, we use Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022) which are effective at making ML models more interpretable. CBMs map a model’s input on to a set of human-understandable concepts, referred to as the “bottleneck” and then use them to make a prediction. However, CBMs require concept annotation for every sample in the training data, which can be costly to obtain. Recently (Yang et al., 2023b; Yan et al., 2023) showed that this limitation of CBMs can be overcome by leveraging the advances in Large Language Models (LLMs) and Vision Language Models (VLMs) that can be prompted to generate concept annotation for the training samples without requiring any task-specific fine-tuning. Our Fig. 1 (block 1) shows the prompt we used to obtain the concept bottleneck using an LLM. Once the bottleneck is formed, we use a VLM (such as CLIP (Radford et al., 2021)) to measure the alignment between the visual features and the concept bottleneck (denoted as concept similarity in Fig. 1 (block 2)). A simple linear concept bottleneck layer is then trained to align the visual and concept features to make predictions on the training samples. We use the average margin of a training sample during the training of this bottleneck layer as our concept-based difficulty score. We use a stratified sampling (Zheng et al., 2022) (block 3) using our score to form the coreset which can then be used to train various downstream models Fig. 1 (block 4). Our method speeds up the computation of the coreset by $15$ times compared to approaches based on the downstream model’s training dynamics across three benchmark datasets. Moreover, compared to coresets found by other training dynamics-free approaches our coresets lead to downstream models with significantly better accuracy (see Sec. 5).
We empirically evaluate the effectiveness of our score on three benchmark datasets: CIFAR-10/100 (Krizhevsky et al., 2009) and Imagenet-1K (Deng et al., 2009). Our results show that downstream models trained on our coresets consistently achieve better accuracy than randomly sampled subsets, especially at high data pruning rates, and achieve performance close to SOTA CS methods. We also show that our approach is effective for the label-free CS problem where the dataset is unlabeled and leads to models with superior performance compared to SOTA label-free CS methods on Imagenet. Since our CS method is model agnostic, we show that our coreset leads to high performance regardless of the architecture of the downstream model, unlike other methods which require training the downstream model on the dataset first to identify the coresets for a new model architecture. Lastly, we show that our concept-based difficulty score is aligned with human intuition, allowing us to interpret why examples are easy/hard, independent of the downstream model. Our main contributions are as follows:
  • We propose a concept-based score for CS that gauges sample difficulty in a model-agnostic way without training a downstream model on the full dataset, speeding up CS by $\approx$15x compared to SOTA methods.
  • Our coresets improve model accuracy by $5$% on average over random subsets at high pruning rates, are competitive to SOTA methods, generalize to various architectures, and are computable without labels.
  • We show that using LLM-generated concepts makes our score human-understandable, enabling a generalizable and data-centric solution for CS.
2 Related Work
Coreset selection (CS): CS improves the efficiency of model training by selecting a subset of influential samples. Various approaches have been proposed to generate such a subset (Guo et al., 2022). A popular approach uses influence functions (Koh & Liang, 2017; Chatterjee & Hadi, 1986; Liu et al., 2021; Schioppa et al., 2022) which measures the influence of a sample by considering the effect of removing it from the model’s training. While effective, these approaches are computationally costly due to their dependence on higher-order derivatives. Another set of approaches selects a subset by either matching the gradients to those computed on the entire dataset (Mirzasoleiman et al., 2020; Killamsetty et al., 2021) or uses training dynamics of a model (Toneva et al., 2018; Pleiss et al., 2020; Lewis & Catlett, 1994; Culotta & McCallum, 2005; Paul et al., 2021) to compute the importance of a sample. However, such approaches require repeated training of the downstream model to produce accurate importance scores. In comparison, our approach avoids using any knowledge of the downstream model for computing the difficulty scores. Finally, using the dataset’s geometric properties via clustering is another popular choice for CS (Sener & Savarese, 2017; Sorscher et al., 2022; Feldman et al., 2020; Feldman & Langberg, 2011; Huang et al., 2019). However, the high computational complexity due to their dependence on pairwise distances between the samples prohibits their use on large datasets.
Concept-based interpretability approaches: Concepts are defined as high-level semantics that refers to the abstract and human-interpretable meanings of the visual data, such as objects, actions, and scenes, as opposed to low-level features like edges or textures (Wu et al., 2016). Concepts have been used in interpretable computer vision to bridge the gap between human understanding and machine perception in various tasks such as image classification. Such interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Concept Bottleneck Models (CBMs) extend interpretable-by-design approaches by using human-understandable attributes as an intermediate layer for predictions, as used in few-shot learning (Lampert et al., 2013) and attribute learning (Xu et al., 2020; Russakovsky & Fei-Fei, 2012). While interpretable, CBMs reliance on costly annotations and lower accuracy compared to end-to-end models limit their usage. Post-hoc Concept Bottleneck Models (PCBMs) address these issues by incorporating static knowledge bases (e.g., ConceptNet (Speer et al., 2017)) and residual connections to boost accuracy (Yuksekgonul et al., 2022). Recently (Yang et al., 2023b; Yan et al., 2023) incorporated LLMs to identify the concept bottleneck to make classification more explainable. We build on this literature and use CBMs for CS. See App. A for other related work.
3 Preliminaries
3.1 Coreset selection (CS) problem formulation
Consider a classification task and data distribution $P$. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the dataset of $n$ training examples sampled i.i.d. from the distribution $P$ where $x_{i}$ denotes the data and $y_{i}\in\mathcal{Y}$ denotes the label from a set of $N$ classes. CS (Coleman et al., 2019; Zheng et al., 2022) aims to find a subset $\mathcal{S}$ of $\mathcal{D}$ consisting of $m\leq n$ samples such that the models trained on $\mathcal{S}$ achieve performance comparable to models trained on $\mathcal{D}$. Formally, the CS problem is
$\min_{\mathcal{S}:|\mathcal{S}|=m}\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{S}})]-\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{D}})],$ (1)
where $\theta_{\mathcal{D}}$ and $\theta_{\mathcal{S}}$ denote the “downstream model” trained on $\mathcal{D}$ and $\mathcal{S}$ (coreset), respectively and $\ell$ is the loss function.
To find this subset $\mathcal{S}$, previous works have proposed scores to gauge a sample’s difficulty for a model, and are later used to form the coreset. Approaches such as max entropy uncertainty sampling (Lewis & Catlett, 1994; Settles, 2012), and least confidence (Culotta & McCallum, 2005) estimate difficulty using the uncertainty of the model’s predictions on a sample. Another set of approaches such as $k$-center greedy (Sener & Savarese, 2017) uses geometric information of the data to filter out redundant samples. Yet, another set of approaches uses information from the training dynamics of the downstream model to estimate the a difficulty score. Scores such as the forgetting score (Toneva et al., 2018) which is computed as the number of times a sample gets misclassified after being correctly classified earlier during model training, the area under the margin (AUM) (Pleiss et al., 2020) which identifies mislabeled/difficult samples, fall in this category. While approaches based on the training dynamics of the downstream model have achieved SOTA results, the requirement of knowledge/training the downstream model or a relatively big proxy model on the entire dataset at least once is inefficient for large datasets/models, even if done only once. This motivates the need of data centric approaches for sample’s difficulty independent of the downstream model.
3.2 Concept bottleneck models (CBMs)
Recent advances in language model-guided CBMs utilize an LLM to obtain concept bottlenecks which are then used to predict the labels. These works rely on a pre-trained multi-modal models (such as CLIP (Radford et al., 2021)) which consists of a visual encoder $\mathcal{V}_{enc}$ and a text encoder $\mathcal{T}_{enc}$ that can map images and text to a $d$-dimensional representation space. Let $C=\{c_{1},c_{2},\cdots,c_{N_{C}}\}$ be the set of $N_{C}$ concepts (bottleneck) generated via a LLM, we can then construct a bottleneck embedding matrix $E_{C}\in\mathcal{R}^{N_{C}\times d}$ such that each row of the matrix is mapping of the concept $c\in C$ after passing it through textual encoder $\mathcal{T}_{enc}$. Based on this, a CBM (Yang et al., 2023b) produces a prediction $h(x)=f(g(\mathcal{V}_{enc}(x),E_{C}))$ for a sample $x$ where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N_{C}}$ computes the similarity of the visual features to each concept in the bottleneck and $f:\mathbb{R}^{d}\rightarrow\Delta$ outputs the probability of each class in the label set $\mathcal{Y}$, where $\Delta$ is a $N$ simplex. We discuss details of $f$ and $g$ in Sec. 4.
4 Methodology
4.1 Generating the concept bottleneck via LLMs
Since obtaining data with concept annotation is costly we use LLMs to generate concept annotation for the samples. However, generating attributes (word-level concepts) for all the samples in the dataset via LLMs is still costly, hence we generate attribute-level concepts only for class label names. This approach was recently shown to be effective at generating the concept bottleneck for interpretable image classification  (Yan et al., 2023; Yang et al., 2023b). In Figure 1 (block 1), we present the prompts provided to the LLMs to extract the concepts for various class label names. The responses of the LLM are then processed to remove formatting errors to obtain the concept sets. Details of our prompt design, robustness check of different prompts, and examples of the generated attributes are mentioned in App. B. Once the per-class concepts are extracted, we select $k$ discriminative concepts per class (concepts that are unique to a class) to form the concept bottleneck. Our final list of $k$ concepts for a class contains its name and $k-1$ attributes generated by the LLM. In Sec. 5.5 we show an ablation study using different methods and LLMs to obtain the concept sets, and the number of concepts $k$.
4.2 Concept-based score for CS
Next, we describe how to use the concept bottleneck to produce a difficulty score for the samples in the dataset. We start by discussing how we learn the functions $f$ and $g$ described in Sec. 3.2 (see Fig. 1(block 2) for an overview of the method). We use the dot product between the visual embeddings of an image $x$ denoted as $\mathcal{V}_{enc}(x)$ and the bottleneck embedding matrix $E_{C}$ to measure the alignment between the visual features and textual features of the concepts (Yang et al., 2023b; Yan et al., 2023). Concretely, the concept similarity scores for a sample $x$ is computed as
$g(x,E_{C}):=\mathcal{V}_{enc}(x)\cdot E_{C}^{\intercal}.$ (2)
To map the concept similarity scores to predictions in the label space $\mathcal{Y}$, we propose to use a linear (concept bottleneck layer) predictor as $f$. Concretely, the function $f$ with parameters $W\in\mathbb{R}^{N\times N_{C}}$ is given by $f(x;W):=g(x,E_{C})\cdot W^{\intercal}$. We learn the parameters $W$ for the function $f$ using
$W^{*}=\arg\min_{W}\frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i};W),y_{i}),$ (3)
where $\ell(f(x;W),y)=-\log(f(x;W)_{y})$ is the cross-entropy loss. We define the concept bottleneck layer’s output, $h(x):=f(g(x,E_{C});W^{*}).$ In practice, we learn the optimal $W$ using mini-batch gradient descent by running the optimization for $T$ epochs. To gauge the difficulty of each training sample we use the area under the margin (AUM) (Pleiss et al., 2020) while solving Eq. 3 which quantifies the data difficulty as the margin of a training sample averaged over $T$ training epochs. Concretely, the margin of a sample $(x,y)$ at a training epoch $t$ is $M^{t}(x,y)=h^{t}_{y}(x)-\max_{y^{\prime}\neq y}h^{t}_{y^{\prime}}(x),$ where $h^{t}_{y^{\prime}}(x)$ is the prediction likelihood of the bottleneck layer at epoch $h^{t}$ for class $y^{\prime}$. Thus, AUM (concept-based score) is computed as
$\mathrm{AUM}(x,y)=\frac{1}{T}\sum_{t=1}^{T}M^{t}(x,y).$ (4)
Recent works (Pleiss et al., 2020; Zheng et al., 2022, 2024) have demonstrated the effectiveness of AUM for gauging the sample’s difficulty for coreset selection. However, (Zheng et al., 2022, 2024) requires computing AUM for the downstream model by training it on the entire dataset once, which is computationally costly. On the other hand, our method integrates AUM with the training of the linear concept bottleneck layer $h$, which is computationally cheaper (training a linear layer takes only $7$ minutes for Imagenet compared to $8$ hours for training a ResNet-34) than training the downstream model ($\theta$). Moreover, since our method does not use the downstream model in the computation of the difficulty score, the coresets are independent of the model architecture, unlike previous training dynamics-based approaches.
Sampling training examples to form a coreset. After obtaining data difficulty scores, a crucial step is choosing the samples to form the coreset. While many previous works (Toneva et al., 2018; Coleman et al., 2019) have reported encouraging results keeping the most challenging samples (for our concept-based score this means samples with the smallest margin), recent works (Zheng et al., 2022; Sorscher et al., 2022) have shown that this could lead to a catastrophic drop in accuracies after training the downstream model on the coreset, especially when the size of the coreset is small. This is mainly due to poor sample coverage and potentially mislabeled data in the datasets. To remedy this, we use Coverage-centric Coreset Selection (CCS) proposed by (Zheng et al., 2022) (see Alg. 1 in App. C.2) which uses a stratified sampling approach and filters out (potentially) mislabeled samples to form the coreset. This technique has been shown to achieve consistently superior results to the random baselines for various coreset sizes.
4.3 Concept-based score for label-free CS
Recently, there has also been an interest (Zheng et al., 2024; Maharana et al., 2023; Griffin et al., 2024) in identifying the representative samples from an unlabeled dataset such that 1) we reduce the samples that need to be labeled by humans and 2) we can improve the efficiency of model training by only training the model on a subset of data. Our concept-based score can also be effectively utilized for this task with a simple modification. Similar to previous works (Maharana et al., 2023; Zheng et al., 2024; Sorscher et al., 2022), we assume that we know the number of classes in the datasets. Additionally, we assume that we also know the names of the classes in the datasets. Previous works have demonstrated that VLMs such as CLIP (Radford et al., 2021) achieve excellent zero-shot performance without requiring fine-tuning on specific datasets. We leverage this capability of CLIP models to obtain pseudo-labels for the images in our unlabeled dataset and use them to obtain our difficulty score for each sample
$\mathrm{AUM}(x,y_{\texttt{pseudo}})=\frac{1}{T}\sum_{t=1}^{T}M^{t}(x,y_{\texttt{pseudo}}),$ (5)
where for an image $x$ in the dataset, $y_{\texttt{pseudo}}=\arg\max_{j\in\mathcal{Y}}\mathcal{V}_{enc}(x)\cdot\mathcal{W}^{\intercal}_{\texttt{zeroshot}}$ where $\mathcal{W}_{\texttt{zeroshot}}\in\mathbb{R}^{N\times d}$ is a matrix with columns defined as $\mathcal{T}_{enc}(s_{j})$ and $s_{j}=$ “a photo of a {$j^{th}$ class name}” for each class $j\in\mathcal{Y}$ (Radford et al., 2021; Wortsman et al., 2022). We use these scores along with CCS to produce the coreset. Similar to (Zheng et al., 2024; Maharana et al., 2023), this coreset is then assumed to be annotated by humans and downstream models ($\theta$) are trained on this annotated coreset.
5 Experiments
Here, we compare the performance of the models trained on coresets generated by our concept-based score to models trained on coresets selected by various SOTA CS methods for the standard and label-free CS problems.
Table 1: Comparison of the model’s (RN-18 for CIFAR10/100 and RN-34 for Imagenet) test accuracy after training on coresets found by various approaches shows that our coresets lead to significantly better performance than Random achieving competitive results compared to the methods using the downstream model’s training dynamics, even for high pruning rates. (Best in each category is highlighted).
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100 Imagenet
$30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$
Needs Training Dynamics Entropy $94.44$ $92.11$ $85.67$ $66.52$ $72.26$ $63.26$ $50.49$ $28.96$ $72.34$ $70.76$ $64.04$ $39.04$
Forgetting $\mathbf{95.40}$ $\mathbf{95.04}$ $92.97$ $85.70$ $\mathbf{77.14}$ $\mathbf{74.45}$ $\mathbf{68.92}$ $\mathbf{55.59}$ $\mathbf{72.60}$ $\mathbf{70.89}$ $66.51$ $52.28$
AUM $95.27$ $94.93$ $\mathbf{93.00}$ $\mathbf{86.08}$ $76.84$ $73.77$ $68.85$ $55.03$ $72.29$ $70.52$ $\mathbf{67.78}$ $\mathbf{57.36}$
Doesn’t Need Training Dynamics Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
RandomFFCV - - - - - - - - $73.37$$\pm 0.08$ $71.71$$\pm 0.10$ $67.85$$\pm 0.04$ $51.29$$\pm 0.20$
Ours $\mathbf{94.77}$$\pm 0.09$ $\mathbf{93.44}$$\pm 0.61$ $\mathbf{91.80}$$\pm 0.21$ $\mathbf{84.63}$$\pm 0.24$ $\mathbf{75.98}$$\pm 0.26$ $\mathbf{72.22}$$\pm 0.22$ $\mathbf{66.53}$$\pm 0.42$ $\mathbf{51.85}$$\pm 0.29$ $\mathbf{73.39}$$\pm 0.12$ $\mathbf{72.34}$$\pm 0.13$ $\mathbf{69.44}$$\pm 0.17$ $\mathbf{55.92}$$\pm 0.02$
Table 2: Comparison of the model’s (RN-18 for CIFAR10/100 and RN-34 for Imagenet) test accuracy after training on coresets, found in a label-free manner, shows that our coresets lead to better performance than Random and those found by Prototypicality and is competitive/better than the coresets found by ELFS (which is dependent on the training dynamics of the downstream model).
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100 Imagenet
$30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$
Needs Training Dynamics ELFS (SwAV) $95.00$ $94.30$ $91.80$ $82.50$ $76.10$ $72.10$ $65.50$ $49.80$ $73.20$ $71.40$ $66.80$ $53.40$
ELFS (DINO) $\mathbf{95.50}$ $\mathbf{95.20}$ $\mathbf{93.20}$ $\mathbf{87.30}$ $\mathbf{76.80}$ $\mathbf{73.60}$ $\mathbf{68.40}$ $\mathbf{54.90}$ $\mathbf{73.50}$ $\mathbf{71.80}$ $\mathbf{67.20}$ $\mathbf{54.90}$
Doesn’t Need Training Dynamics Prototypicality $94.70$ $92.90$ $90.10$ $70.90$ $74.50$ $69.80$ $61.10$ $32.10$ $70.90$ $60.80$ $54.60$ $30.60$
Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
RandomFFCV - - - - - - - - $73.37$$\pm 0.08$ $71.71$$\pm 0.10$ $67.85$$\pm 0.04$ $51.29$$\pm 0.20$
Ours-LF $\mathbf{94.81}$$\pm 0.14$ $\mathbf{93.93}$$\pm 0.13$ $\mathbf{91.75}$$\pm 0.34$ $\mathbf{84.02}$$\pm 0.44$ $\mathbf{74.67}$$\pm 0.23$ $\mathbf{72.07}$$\pm 0.58$ $\mathbf{65.50}$$\pm 0.17$ $\mathbf{49.91}$$\pm 0.96$ $\mathbf{73.61}$$\pm 0.08$ $\mathbf{71.99}$$\pm 0.05$ $\mathbf{68.42}$$\pm 0.21$ $\mathbf{53.21}$$\pm 0.06$
Datasets, models, and training: We focus on CS for classification tasks on three benchmark datasets. Specifically, we use CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and Imagenet-1K (Deng et al., 2009) datasets consisting of $50000$, $50000$, and $1.28$ million samples spread across $10$, $100$, and $1000$ classes, respectively. For CIFAR-10/CIFAR-100, we train a ResNet-18 model and for Imagenet we train ResNet-18, ResNet-34, and ResNet-50 models on the coresets for all pruning ratios. To accelerate training on Imagenet, we utilize the training code based on FFCV (Leclerc et al., 2023). We run CS for three trials with different random seeds for all experiments and report the average of these runs in our tables. We report the results for different pruning rates where a pruning rate of $90$% refers to removing $90$% of the samples from the original training dataset.
For generating the concept annotation we use a recently proposed open source model LLaVA (Liu et al., 2023b, a). Extracting the concepts for each class using this model takes a mere 3 seconds per prompt. In Sec. 5.5, we present an ablation study using concepts extracted from different models/methods. For computing the concept similarity scores between the visual and concept bottleneck features we used the CLIP (Radford et al., 2021) model following the previous works (Yun et al., 2022; Yang et al., 2023b; Yan et al., 2023) which showed its effectiveness for this task. Specifically, we used the ViT B-32 CLIP model (Radford et al., 2021). For computing the pseudo-labels in Sec. 4.3 we used a ViT L-14 CLIP model trained on the DataComp-1B dataset (Ilharco et al., 2021). The accuracies of the models trained on the entire training set are $95.44$% and $78.74$% for ResNet(RN)-18 on CIFAR-10/100 and $72.4$% for RN-18, $75$% for RN-34, and $78.4$% for RN-50 on Imagenet. Further experimental details are mentioned in App. C.
CS baselines and methods: We compare our method against various baselines and SOTA methods proposed by previous works. 1) Random: Uniformly select samples from the datasets to form the coreset. RandomFFCV denotes the performance of the models trained on random subsets of Imagenet using the training code based on FFCV (Leclerc et al., 2023). 2) Entropy (Coleman et al., 2019): Selects samples based on entropy which is computed as the uncertainty in the model’s prediction on a sample. 3) Forgetting Score (Toneva et al., 2018): Selects samples based on the forgetting score which is computed as the number of times an example is misclassified after being correctly classified earlier during training of the downstream model. A higher forgetting score indicates a more challenging sample. 4) AUM (Pleiss et al., 2020): Selects samples based on their average margin during training of the downstream model i.e., the difference between the target class and the next highest class across the training epochs. Lower AUM indicates a more challenging sample. For the forgetting score, AUM, and our method, we use CCS (Zheng et al., 2022) for sampling to form the coreset whereas for entropy we select the samples with the highest entropy in the coreset as done in previous works (Coleman et al., 2019; Zheng et al., 2022).
For label-free CS, we use 1) Prototypicality (Sorscher et al., 2022): which first performs k-means clustering in the embedding space of SwAV (Caron et al., 2020) model and ranks samples based on their Euclidean distance to the cluster centers. Samples further away from the cluster center are then used to form the coreset. 2) ELFS (Zheng et al., 2024): estimates the pseudo-labels of the unlabeled samples using a deep clustering approach (using the embedding space of SwAV (Caron et al., 2020) and DINO (Caron et al., 2021)) and forms the coreset using the training dynamics of the downstream model trained on the pseudo-labeled data.
Crucially, SOTA methods such as forgetting score, AUM, and ELFS require training the downstream model on the entire dataset (with true/pseudo labels) first, for CS, unlike our method which is independent of the downstream model. While the Random and the Prototypicality don’t require the downstream model for CS, we show in the following sections that our results are significantly better than these.
5.1 Evaluating performance on CS
Table 1 shows the accuracy of models trained on coresets found by various approaches on the test sets of the three datasets for the standard CS problem (where the dataset is labeled). We find that coresets found by our approach lead to significantly better performance, even at higher pruning rates, compared to the random subsets. Moreover, our method which does not use any information about the training dynamics of the downstream models provides competitive performance to coresets found by the SOTA approaches based on forgetting score and AUM, and even outperforms them on Imagenet for smaller pruning rates.
5.2 Evaluating performance on label-free CS
Table 2 shows the accuracy of models trained on coresets when the training set lacks labels (we report the numbers presented by (Zheng et al., 2024) for all the methods and baselines). The results show that the random subsets are a competitive baseline and even it outperforms Prototypicality (Sorscher et al., 2022). Our results also show that our coresets outperform the random subsets for all pruning rates with the improvements being the most significant at higher pruning rates. Compared to ELFS (Zheng et al., 2024), our method provides competitive performance and even surpasses it for lower pruning rates on Imagenet, without using any information about the downstream model’s architecture or its training dynamics. Thus, our method is effective for this problem with a simple modification of utilizing the zero-shot predictions from CLIP as pseudo-labels.
Figure 2: Visualizing samples according to our concept-based score for a subset of classes in CIFAR-10/100 showing that easy (challenging) samples are aligned (unaligned) with their assigned label. Image-level concepts (in boxes) extracted via LLaVA confirm that easy (challenging) examples are aligned (unaligned) with concepts of their labels, explaining the reason for a high (low) concept-based score.
5.3 Evaluating the transferability and efficiency of CS
Here we evaluate the performance of training downstream models with three different architectures to highlight the effectiveness of our approach for model-agnostic CS. Table 3 shows superior performance than random for ResNet-18, ResNet-34, and ResNet-50 models trained on our coresets for standard and label-free CS for various pruning rates.
Next, we compare the efficiency of our approach at finding coresets compared to approaches relying on training the downstream model on the entire dataset to obtain the training dynamics-based score. Using two A-100 GPUs, our approach can find the coreset in approximately $30$ minutes for the Imagenet dataset giving a $15$x speed up over training dynamics-based approaches. We obtain a similar speed up for CIFAR-10/100 where our method finds the coreset in less than $2$ minutes. Moreover, since our method is agnostic to the architecture of the downstream model we do not need to repeat the CS step for different architectures, unlike other methods which need to train the downstream model with new model architecture to find the best coreset for it.
Table 3: Superior performance of downstream models with different architectures trained on our coresets for Imagenet compared to Random for standard (Ours) and label-free (Ours-LF) CS highlight our method’s effectiveness for downstream model-agnostic CS.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
RN-18 Random $71.15_{\tiny\pm 0.23}$ $68.48_{\tiny\pm 0.10}$ $63.15_{\tiny\pm 0.19}$ $44.96_{\tiny\pm 0.50}$
Ours-LF $71.21_{\tiny\pm 0.09}$ $68.77_{\tiny\pm 0.13}$ $63.76_{\tiny\pm 0.09}$ $47.50_{\tiny\pm 0.10}$
Ours $70.94_{\tiny\pm 0.19}$ $69.30_{\tiny\pm 0.84}$ $65.16_{\tiny\pm 0.04}$ $49.57_{\tiny\pm 0.16}$
RN-34 Random $73.37_{\tiny\pm 0.08}$ $71.71_{\tiny\pm 0.10}$ $67.85_{\tiny\pm 0.04}$ $51.29_{\tiny\pm 0.20}$
Ours-LF $73.61_{\tiny\pm 0.08}$ $71.99_{\tiny\pm 0.05}$ $68.42_{\tiny\pm 0.21}$ $53.52_{\tiny\pm 0.06}$
Ours $73.39_{\tiny\pm 0.12}$ $72.34_{\tiny\pm 0.13}$ $69.44_{\tiny\pm 0.17}$ $55.92_{\tiny\pm 0.02}$
RN-50 Random $76.06_{\tiny\pm 0.11}$ $74.44_{\tiny\pm 0.04}$ $70.50_{\tiny\pm 0.02}$ $53.56_{\tiny\pm 0.13}$
Ours-LF $76.54_{\tiny\pm 0.08}$ $74.84_{\tiny\pm 0.03}$ $71.10_{\tiny\pm 0.09}$ $55.22_{\tiny\pm 0.92}$
Ours $76.29_{\tiny\pm 0.10}$ $75.12_{\tiny\pm 0.03}$ $72.05_{\tiny\pm 0.09}$ $58.26_{\tiny\pm 0.46}$
5.4 Visualizing samples with their concept-based scores
Here we show the advantage of our concept-based score for assessing the sample’s difficulty and how using concepts aids its interpretability. We start by visualizing the easiest and the most challenging images (per class) for CIFAR-10/100. In Fig. 2, the top row shows the images with the highest concept-based scores (easiest) and the bottom shows the images with the lowest scores (challenging) for a subset of classes in CIFAR-10/100. As observed the easiest images are typical images associated with the label where as the challenging images are confusing (and even potentially mislabeled) as they look like images from a different class. For example, some challenging images in the class “boy” from CIFAR-100 are actually images of a baby which is also a class in CIFAR-100. Similarly, some challenging images from the class “cat” in CIFAR-10 look like images of a dog. More examples of such images are presented in Fig. 3 in App. C.1. Since the challenging examples are confusing for humans too, any ML model will find them hard to learn as well. This shows that our score is well aligned with the human-perceived difficulty of the samples.
Next, we demonstrate why certain samples get low/high concept-based scores in our approach by extracting attributes specific to these images using LLaVA (note that these attributes are different from the per-class concepts used in the concept bottleneck). To generate these, we prompt LLavA to produce concepts using both the sample image and its class label (see image-level concept extraction in App. B). These image-level attributes are shown in the boxes in Fig. 2. As observed in Fig. 2, image-level attributes provided by LLaVA are related to the class label for easy images whereas they are unrelated for challenging images. For example, attributes provided by LLaVA for the challenging images of “airplane” align more with those of a ship (both of which are classes in CIFAR-10), and concepts provided for challenging images of “bridges” align more with those of castles (both of which are classes in CIFAR-100). Since our score also assigns a small value for these images, our concept-based score in Eq. 4 can correctly capture when the visual information in the sample is not aligned with the associated label of the sample and vice-versa. Thus, explaining why certain examples should be included/excluded from the coreset in a human-understandable way and at a dataset level independent of the downstream model.
5.5 Ablation studies
Here, we present an ablation to study the effect of different methods for concept generation and the effect of keeping different numbers of concepts per class ($k$) in the bottleneck.
Comparison of different techniques to generate concepts via LLaVA. Table 4 shows how the performance of models trained on our coresets change when different methodologies are used to generate the concept sets. Since our method uses LLaVA (Liu et al., 2023b), which is a vision language model, we compare the performance of models trained on the random subsets and coresets obtained using class-wise concepts (only textual information) and concepts extracted using both visual and textual information. For concepts generated using only textual information, we consider two alternatives, namely class-wise attributes (CW-A) and class-wise descriptions (CW-D). While CW-A considers concepts formed by a single or a few words, CW-D consists of longer, more descriptive concepts (eg., a descriptive concept for the class butterfly is “a beautiful insect with colorful wings”). For CW-D, we use a subset of $k$ concepts provided by (Yang et al., 2023b), generated via the GPT-3 model. Our results show that CW-A performs better than CW-D for both the standard and label-free CS problems. Thus, we use attribute-level concepts in our work.
Table 4: Coreset performance on CIFAR-100 for $90\%$ pruning rate for concepts generated using different methods. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\tiny\pm 1.58}$ at this rate. (Class-wise attributes approach is used for in rest of the paper.)
Concept generation methods
Prompt uses only class label Prompt uses class label + image(s)
class-wise attributes class-wise descriptions class-wise one-shot image attributes image-wise attributes
Ours $51.85_{\tiny\pm 0.51}$ $51.05_{\tiny\pm 0.71}$ $51.68_{\tiny\pm 0.45}$ $52.47_{\tiny\pm 0.39}$
Ours-LF $49.91_{\tiny\pm 0.96}$ $49.85_{\tiny\pm 0.83}$ $50.22_{\tiny\pm 0.16}$ $51.05_{\tiny\pm 0.93}$
Next, for generating concepts using both visual and textual information, we consider two alternatives. The first is a class-wise one-shot image attribute approach where we first cluster all images of a class in the embedding space of the CLIP’s visual encoder and identify the image whose embedding is the closest to the cluster center (for the label-free setting we use the pseudo-labels of the images during clustering ), then we prompt LLaVA to generate attributes using this single image and the class name. Once generated we use $k$ discriminative concepts to form the bottleneck. The second is image-wise attribute approach, where we use each image in the training set and prompt LLaVA to generate per image attributes describing the image. Once generated we sort the concepts based on their frequency of occurrence in a class and use the most frequently occurring discriminative concepts to form the bottleneck. While the image-level concepts lead to the best coresets, it is slow and costly to prompt LLaVA to generate attributes for all the images in a large dataset such as Imagenet. For CIFAR-100, this process took about nine hours to complete (in comparison CW-A can be extracted in 5 minutes for CIFAR-100, without parallel computation) which is very costly compared to the small performance gains it provides over other approaches. Lastly, while the one-shot approach is better than CW-A in most cases, the additional step of clustering can be costly for larger datasets such as Imagenet. Based on these results, we used CW-A for concept generation using LLaVA.
Effect of $k$. In Table 5, we show how the number of concepts extracted per class label affects the selection of coresets. Once the list of attribute-level concepts is generated by LLaVA, we can select $k$ concepts per class either randomly or choose concepts unique to a class (discriminative). Our results show that using even $k$=$1$ is sufficient to surpass the performance using a random subset (Random baseline). This performance increases when we keep discriminative concepts in our concept bottleneck, with $k=5$ achieving the best results. While the size of the concept bottleneck need not be very large to find good coresets, it is helpful to take a sample’s visual similarity with a set of concepts rather than a single concept per class. Thus, we selected $5$ discriminative concepts per class to form the concept bottleneck.
Table 5: Effect of the number of concepts $k$ (per class) and the method of selecting $k$ attribute-level concepts from the concepts generated by LLaVA vs. the accuracy of models trained on the coresets of CIFAR-100 at $90\%$ pruning rate. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\tiny\pm 1.58}$ at this rate.
Concept selection $k$=1 $k$=5 $k$=10
Random concepts $48.78_{\tiny\pm 0.96}$ $50.15_{\tiny\pm 1.64}$ $50.39_{\tiny\pm 0.72}$
Discriminative $51.42_{\tiny\pm 0.18}$ $\mathbf{51.58}_{\tiny\pm 0.51}$ $51.22_{\tiny\pm 0.72}$
6 Conclusion
CS finds representative samples from the training data, training models on which leads to models with accuracy similar to the models trained on the entire dataset. In this work, we proposed a score based on concept bottlenecks that allow us to gauge the difficulty of a sample in terms of human-understandable concepts and is independent of the downstream model. Our experiments show that training downstream models on coresets selected using our score and a stratified sampling approach leads to better performance than random subsets and achieves accuracy similar to or better than the SOTA approaches based on training dynamics of the downstream model, for both the standard and label-free CS problem. Moreover, our score provides an intuitive explanation of a sample’s difficulty independent of the downstream model and leads to coresets that achieve high accuracy for various model architectures.
Limitations: In our work, the concept extraction from LLM is treated as a pre-processing step which is independent of the data difficulty score computation and coreset selection. However, LLMs can produce very noisy and non-discriminative concepts for various classes (eg., a concept “blue” can be associated with several classes in the Imagenet dataset), leading to poor concept-similarity scores. Moreover, while class-wise concept extraction is efficient, image-level concepts could be much more informative for a sample’s difficulty. Thus, improving the efficiency of concept extraction and tuning the prompt for LLMs/VLMs to incorporate the feedback from score computation or CS can help generate better concept bottlenecks leading to a better estimate of a sample’s difficulty. While these are important research directions we leave them for future work.
Appendix
We present additional related work in Appendix A. Then we describe the details of our methodology for extracting the concepts from LLaVA in Appendix B and present additional experiments and implementation details of our experiments in Appendix C including the algorithm for stratified sampling used in our work in Appendix C.2.
Appendix A Additional related work on concept-based interpretability
Interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Post-hoc methods include Gradient-weighted Class Activation Mapping approaches (Bau et al., 2017; Selvaraju et al., 2017; Mu & Andreas, 2020; Hernandez et al., 2021) that trace network gradients to identify the input areas that guide predictions and Explanation Generation methods (Singh et al., 2023; Nishida et al., 2022; Kim et al., 2018; Hendricks et al., 2016) that require models to produce explanations for visual tasks by conditioning their predictions on captioning models or incorporating visual evidence to ground explanations (Hendricks et al., 2018; Park et al., 2018). Interpretable-by-design methods, such as Prototype methods, optimize a metric space where classifications are based on distances to class prototypes, identifying important input regions but often obscuring their semantic content (Nauta et al., 2021; Chen et al., 2019; Snell et al., 2017; Satorras & Estrach, 2018; Vinyals et al., 2016).
Concept Bottleneck Models (CBMs) are a part of interpretable-by-design approaches that use human-understandable attributes as an intermediate layer for predictions. A recent advancement, Computational Derivation Learning (CompDL), utilizes a CBM architecture by applying a linear layer over CLIP scores between expert-designed concepts and images, improving evaluation of how well CLIP grounds concepts (Yun et al., 2022). Post-hoc Concept Bottleneck Models (PCBMs) were recently proposed to ease the requirement of CBMs to rely on costly concept annotations and improve their accuracy compared to end-to-end models. However, PCBMs are limited by the coverage of knowledge bases, making them unsuitable for large-scale or domain-specific tasks or fine-grained classification, and their residual predictors can undermine interpretability by blending CBMs with end-to-end models.
High-level semantic-driven descriptions are also used to guide data augmentation to build an informative set (Wickramanayake et al., 2021) to make model training efficient with a good enough training set. Prior works use external knowledge bases to obtain these textual semantic concepts to guide vision models (Bujwid & Sullivan, 2021; Kil & Chao, 2021; Roth et al., 2022; Shen et al., 2022). Thus, the use of concepts has been shown to improve interpretability in various domains. However, to the best of our knowledge, we are the first ones to propose a concept-based score for the CS problem and show its competitiveness to SOTA model training dynamics-dependent approaches.
Appendix B Details for concept set generation
[…]
Alternative VLMs for Concept Generation: We leverage LLaVA as our choice of VLM for concept generation, however in Table 4, we also compare against concepts extracted from GPT (column 2) (Yan et al., 2023). We see comparable performance with those extracted from LLaVA. Moreover, LLaVA is an open-source model whereas GPT prompting is not. We also experimented with retrieving concepts via SpLiCE (Bhalla et al., 2024). However, a major limitation of SpLiCE is that similar to image-level concepts it is a costly approach. SpLiCE uses a linear optimization for sparse concept decomposition and can take up to 3 hours for $50,000$ images which is slower than generating class-level concepts from LLaVA.
Appendix C Additional experiments and implementation details
C.1 Visualizing easy/challenging samples based on concept-based score
Similar to Fig 2 in Sec. 5.4 of the main paper, we visualize easy and challenging examples in Fig. 3 for CIFAR-10 and subset of classes from CIFAR-100. As observed the easy images (the ones that get high scores in our approach) are more canonical images of the class labels whereas the challenging ones are images that can potentially be assigned another class in the same dataset or are mislabeled in the dataset. The clear distinctions between these images show that our concept-based score aligns well with human intuition on the difficulty of the samples.
C.2 Algorithm for stratified sampling using CCS (Zheng et al., 2022)
Here we present the algorithm for sampling the training examples to form the coreset based on the coverage-based selection methodology proposed by (Zheng et al., 2022). A crucial component of the algorithm is the cutoff rate $\beta$ which controls how many challenging samples should be removed from consideration when selecting the coreset. This is done to eliminate misclassified samples from the dataset since they can hurt the performance of the model trained on coreset, especially at high pruning rates. Previous works (Zheng et al., 2022, 2024) ablate the values of this cutoff ratio by training the downstream model on a range of values. In our work, we simply use the values proposed by the previous works and find that they work well for our score as well. The cutoff rates $\beta$ for different pruning rates $\alpha$ are as follows ($\alpha$, $\beta$). For CIFAR-10: (30%, 0), (50%, 0), (70%, 10%), (90%, 30%), for CIFAR-100: (30%, 10%), (50%, 20%), (70%, 20%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%). We used CCS for label-free CS as well and the cutoff rates used were for CIFAR-10: (30%, 0), (50%, 0), (70%, 20%), (90%, 40%), for CIFAR-100: (30%, 0), (50%, 20%), (70%, 40%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%).
Algorithm 1 Coverage-centric Coreset Selection (CCS) (Zheng et al., 2022)
Input: Dataset with difficulty scores: $\mathbb{D}=\{(x,y,s)\}_{i=1}^{n}$, pruning ratio: $\alpha$, cutoff rate: $\beta$, number of bins: $b$.
Output: Coreset: $\mathcal{S}$
  # Prune hardest examples
[…]
  $A_{1},A_{2},\cdots,A_{b}\leftarrow$ Split scores in $\mathbb{D}^{\prime}$ into $b$ bins.
  $\mathcal{B}\leftarrow\{B_{i}:B_{i}\;\mathrm{consists\;of\;samples\;with\;scores\;in\;}A_{i}\;for\;i=1,\cdots,b\}$.
  # Define the size of the coreset
  $m\leftarrow n\times\alpha$.
  while $\mathcal{B}\neq\varnothing$ do
     # Select the bin with the fewest examples
     $B_{min}\leftarrow\arg\min_{B\in\mathcal{B}}|B|$.
     # Compute the budgets for this bin
     $m_{B}\leftarrow\min\{|B|,\left\lfloor\frac{m}{|\mathcal{B}|}\right\rfloor\}$.
     $\mathcal{S}_{B}\leftarrow$ randomly sample $m_{B}$ samples from $B_{min}$.
     $\mathcal{C}\leftarrow\mathcal{C}\bigcup\mathcal{S}_{B}.$
[…]
     $m\leftarrow m-m_{B}.$
  end while
  return $\mathcal{C}$.
Table 6: Performance of concept-based coreset selection for emotion recognition task on Affectnet dataset. Coresets selected by our approach results in superior performance (higher F1 score) compared to Random for standard CS
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
EfficientNet Random $0.607_{\tiny\pm 0.006}$ $0.573_{\tiny\pm 0.025}$ $0.517_{\tiny\pm 0.006}$ $0.347_{\tiny\pm 0.055}$
Ours $0.603_{\tiny\pm 0.006}$ $0.577_{\tiny\pm 0.021}$ $0.537_{\tiny\pm 0.015}$ $0.450_{\tiny\pm 0.035}$
Table 7: Performance of concept-based coreset selection for emotion recognition task on BloodMNIST dataset. Coresets selected by our approach results in superior accuracy compared to Random for standard CS; highlighting our method’s effectiveness across various tasks.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
ResNet-18 Random $94.99_{\tiny\pm 0.27}$ $94.79_{\tiny\pm 0.18}$ $91.71_{\tiny\pm 0.29}$ $86.50_{\tiny\pm 0.66}$
Ours $94.64_{\tiny\pm 0.60}$ $94.06_{\tiny\pm 0.10}$ $92.26_{\tiny\pm 0.19}$ $87.44_{\tiny\pm 1.46}$
C.3 Concept-based CS for emotion recognition and biomedical image recognition
To validate the generalizability of our concept-based coreset selection method beyond image recognition tasks, we apply our concept-based CS approach to the task of emotion recognition and biomedical image recognition.
For emotion recognition, we use the Affectnet dataset (Mollahosseini et al., 2017) for our experiments. AffectNet is a large-scale facial expression dataset designed for training and evaluating affective computing models (Wang et al., 2022). It contains facial images collected from the internet using web search queries for emotion-related keywords in multiple languages. Each image is manually annotated for eight discrete emotion categories: neutral, happiness, sadness, surprise, fear, disgust, anger, contempt. For our experiments, we utilize an openly available version of this dataset  [^1] https://www.kaggle.com/datasets/noamsegal/affectnet-training-data, containing roughly $16000$ training and $14000$ testing samples.
According to our approach we first use LLaVA to extract concepts for the 8 emotion classes, using the following prompt, “What are the facial features that distinguish emotion class name from other emotion types. Focus on changes in eyes, nose, lips, eyebrows, mouth. Give the output separated by commas in a line.”. We get $5-10$ distinctive facial feature concepts for every emotion, for instance for emotion class happy, we get the following concepts, “wide open eyes”, “sparking eyes”, “smiling lips”, “open mouth”, “raised eyebrows”, “flushed cheeks”, “teeth barred”.
To test coreset performance, we use the EfficientNet model (Tan & Le, 2019) and report F1 scores for our coresets in Table 6. When compared against randomly selected coresets for the various pruning ratios, coresets selected via our concept-based approach achieve better performance.
For biomedical image recognition, we use the BloodMNIST (Acevedo et al., 2020) dataset from the MedMNIST (Yang et al., 2021, 2023a) which comprises of images of normal blood cells, captured from individuals without infection, hematologic or oncologic disease and free of any pharmacologic treatment at the moment of blood collection. It consists of a total of $17,092$ images and is organized into 8 classes (basophil,eosinophil,erythroblast,immature granulocytes(myelocytes, metamyelocytes and promyelocytes),lymphocyte,monocyte,neutrophil,platelet).
For this dataset, we first extract concepts for the 8 blood cell types via GPT using the following prompt, “What are the features that can distinguish blood cell class name from rest of the blood cell types on their size, shape, nucleus appearance, and the presence of granules in their cytoplasm”. We obtain 10 concepts for every blood cell type, for instance, for platelets, we get the following concepts, “Smallest blood component”, “No nucleus”, “Granules present”, “Irregular shape”, “Cytoplasmic fragments”, “Variable granule distribution”, “Oval to round shape”, “Small dense granules”, “Lacks chromatin”, “Compact cytoplasmic body”.
To test the coreset performance, we use a ResNet-18 model and report accuracy of our coresets in Table 7. Similar to other results, our method achieves better performance than randomly selected coresets for higher pruning rates and is competitive at lower pruning ratios. This is attributes to the difficulty of calculating concept similarity is the representation space of the CLIP model which is potentially unaware of the terminology used in the medical domain. While replacing CLIP with a VLM that is trained on medical domain can boost the performance of our method, our results highlight that even with access to such a model our approach is able to find better coresets than random subsets.
Our results on these two tasks highlight the versatility of our approach for model-agnostic approach to coreset selection, which is able to find coreset without requiring training the the downstream models on the entire dataset even once.
C.4 Experimental details
For generating the importance score we pre-compute the concept similarity scores for the entire dataset and then train the concept-bottleneck layer (in block 2 of Fig. 1) for $100$ epochs across all experiments. This training only requires 800 seconds for Imagenet which is significantly more efficient than training the ResNet-34 model on Imagenet (requires roughly 8 hours without FFCV and about 4 hours with FFCV on two A-100 GPUs).
After the coresets are selected, we use the setting and code from (Zheng et al., 2022) for training a ResNet-18 model for $40000$ iterations with a batch size of $256$ on the coresets for all pruning ratios for CIFAR-10/CIFAR-100. For Imagenet, we train ResNet-18, ResNet-34, and ResNet-50 models for $100$ epochs on the coresets identified by our method using the training code based on FFCV (Leclerc et al., 2023).
The performance of the label-free CS is dependent on the quality of the pseudo-labels. Compared to the clustering-based approach used by ELFS (Zheng et al., 2024), our approach of using the zero-shot classification ability of CLIP models yields significantly better pseudo-label quality along with being simpler and more efficient to compute. Specifically, for CIFAR-10/100, pseudo-labels of the training set are computed using the CLIP L-14 model trained on the DataComp-1B dataset (Ilharco et al., 2021) yields an accuracy of $98.52$% and $87.28$% whereas for Imagenet it achieves an accuracy of $79.47$% which are better than the best pseudo-label accuracy obtained by the clustering approach in ELFS ($92.5$% and $66.3$% on CIFAR-10/100 and $58.8$% on Imagenet).
(a) Easy images from CIFAR-10
(b) Challenging images from CIFAR-10
(c) Easy images from CIFAR-100
(d) Challenging images from CIFAR-100
Figure 3: Class-wise easy and challenging images for the 10 classes (airplane, car, bird, cat, deer, dog, frog, horse, ship, truck) in CIFAR-10 and for a subset of 10 classes (boy, bridge, camel, cloud, crab, kangaroo, lamp, rose, tiger, train) from CIFAR-100. Similar to the results in Fig. 2, easy images (a,c) are more canonical images associated with the class labels whereas challenging images (b,d) are images that are confused between two or more classes in the dataset.