### v2

# Coreset Selection via LLM-based Concept Bottlenecks

Abstract
Coreset Selection (CS) aims to identify a subset of the training dataset that achieves model performance comparable to using the entire dataset. Many state-of-the-art CS methods select coresets using scores whose computation requires training the downstream model on the entire dataset first and recording changes in the model’s behavior on samples as it trains (training dynamics). These scores are inefficient to compute and hard to interpret, as they do not indicate whether a sample is difficult to learn in general or only for a specific downstream model. Our work addresses these challenges by proposing a score that computes a sample’s difficulty using human-understandable textual attributes (concepts) independent of any downstream model. Specifically, we measure the alignment between a sample’s visual features and concept bottlenecks, derived via large language models, by training a linear concept bottleneck layer and computing the sample’s difficulty score using it. We then use stratified sampling based on this score to generate a coreset of the dataset. Crucially, our score is efficiently computable without training the downstream model on the full dataset even once, leads to high-performing coresets for various downstream models, and is computable even for an unlabeled dataset. Through experiments on CIFAR-10/100, and ImageNet-1K, we show that our coresets outperform random subsets, even at high pruning rates, and achieve model performance comparable to or better than coresets found by training dynamics-based methods.
1 Introduction
Machine learning (ML) pipelines are increasingly demanding more data and compute touvron2023llama ; achiam2023gpt to achieve improved performance on various tasks. While in line with empirical neural scaling laws kaplan2020scaling ; hestness2017deep ; henighan2020scaling ; rosenfeld2019constructive where a model’s performance improves with increasing model and training data size, these improvements come at an unsustainable cost of compute/energy. However, recently sorscher2022beyond ; li2024datacomp demonstrated that data pruning plays a crucial role in enabling an exponential reduction in test error with increasing dataset size, underscoring the importance of data quality over quantity.
Figure 1: Computation of training dynamics-based scores require training the original or a big proxy model on the full dataset, to identify the coreset, making them computationally expensive.
Coreset Selection (CS) mirzasoleiman2020coresets ; guo2022deepcore ; paul2021deep ; xia2022moderate ; maharana2023d2 ; zheng2022coverage ; choi2024bws is a data pruning technique that improves the efficiency of model training by pruning a large dataset and retaining only a small subset of representative samples. Most CS methods work by first using a score to estimate the difficulty/importance of every training sample and then using a sampling strategy that forms the coreset. Many state-of-the-art (SOTA) CS methods use the downstream model’s training dynamics — changes in the model’s behavior on a sample over epochs during training, to generate an importance score for each sample. While this enables a good estimation of the data importance, it requires training the downstream model on the entire dataset at least once, which can be costly when training a large model on a large dataset (see Fig. 1). While coleman2019selection showed that a coreset selected using training dynamics of a mid-sized proxy model (eg, ResNet-18) is effective for a larger downstream model (eg, ResNet-50), training even such a model may not be feasible for large datasets. Moreover, since these scores are dependent on training dynamics of a particular downstream model, they are hard to interpret as they do not inform us about the sample’s importance for another downstream model (without training it first). Thus, in our work we tackle the following question: “How to efficiently estimate the importance of training samples for CS, in a way that is independent of the downstream model and avoids training that model on the full dataset?
Figure 2: Overview of our approach: We start by prompting an LLM to generate concept annotation for $N$ class label names in the dataset and select $k$ most discriminative attributes (per class) to form the concept bottleneck, which are passed through a text encoder ($\mathcal{T}_{enc}$) to obtain the bottleneck embedding matrix ($E_{C}$). The visual information of a training sample $x$ extracted via the visual encoder ($\mathcal{V}_{enc}$) is then aligned to $E_{C}$ using a linear concept bottleneck layer ($h$) trained for $T$ epochs. Our difficulty score for a sample $x$ is then computed as the average margin (i.e., the difference between the likelihoods of the correct and other classes) over $T$ epochs. Finally, the coreset is selected via stratified sampling and is used to train downstream models.
To address this, we use Concept Bottleneck Models (CBMs) koh2020concept ; yuksekgonul2022post , which work by mapping a model’s input onto a set of human-understandable concepts, referred to as the “bottleneck” and use them to make a prediction. However, CBMs require concept annotation for every sample in the dataset, which can be costly to obtain. Recently yang2023language ; yan2023learning showed that off-the-shelf Large Language Models (LLMs) and Vision Language Models (VLMs) can be prompted to obtain concept annotations for training samples without requiring any task-specific fine-tuning (see Fig. 2 (block 1)). Once the bottleneck is formed, we use a VLM (e.g., CLIP radford2021learning ) to measure the alignment between the visual features and the concept bottleneck (denoted as concept similarity in Fig. 2 (block 2)). A linear concept bottleneck layer is then trained to align the visual and concept features while classifying samples. We then use the average margin of a sample while training the bottleneck layer as our concept-based importance score. Finally, we form the coreset using stratified sampling zheng2022coverage (block 3) based on our score, which can then be used for training downstream models (Fig. 2 (block 4)).
We empirically evaluate the effectiveness of our score on three benchmark datasets: CIFAR-10/100 krizhevsky2009learning and Imagenet-1K deng2009imagenet . Our results show that downstream models trained on our coresets consistently achieve better accuracy than randomly sampled subsets, especially at high data pruning rates, and achieve performance close to SOTA CS methods. We also show that our approach is effective for the label-free CS problem where the dataset is unlabeled and leads to models with superior performance compared to SOTA label-free CS methods, especially on Imagenet. Since our CS method is independent of the downstream model, we show that our coreset leads to high performance regardless of the architecture of the downstream model. Moreover, our method speeds up the computation of the coreset by $\approx 8$ times compared to approaches based on the training dynamics of the downstream model. Lastly, we show that our concept-based difficulty score provides an intuitive explanation of why examples are easy/hard, independent of the downstream model. Our main contributions are summarized below:
  • We propose a concept-based score that efficiently computes a training data sample’s importance for CS without training a downstream model on the full dataset, even once.
  • Our coresets improve accuracy by $\approx 5$% over random subsets at high pruning rates, are competitive to SOTA methods, transfer to various architectures, and are computable without labels.
  • We show that using CBMs with LLM-generated concepts makes our score interpretable, enabling a data-centric solution for identifying coresets.
2 Related Work
Coreset selection (CS). CS improves the efficiency of model training by selecting a subset of influential samples. Various approaches have been proposed to generate such a subset guo2022deepcore . A popular approach uses influence functions koh2017understanding ; chatterjee1986influential ; liu2021influence ; schioppa2022scaling which measures the influence of a sample by considering the effect of removing it from the model’s training. While effective, these approaches are computationally costly due to their dependence on higher-order derivatives. Approaches that use the dataset’s geometric properties such as are another popular choice for CS sener2017active ; sorscher2022beyond ; feldman2020turning ; feldman2011unified ; huang2019coresets . However, the high computational complexity due to their dependence on pairwise distances between the samples prohibits their use on large datasets. Another set of approaches select a subset by either matching the gradients to those computed on the entire dataset mirzasoleiman2020coresets ; killamsetty2021grad or use training dynamics of a model toneva2018empirical ; pleiss2020identifying ; lewis1994heterogeneous ; culotta2005reducing ; paul2021deep to compute the importance of a sample. However, such approaches require repeated training of the downstream model to produce accurate importance scores. In comparison, our approach avoids using any knowledge of the downstream model for computing the difficulty scores. Unlike our work which focuses on data pruning, a different line of work killamsetty2023milo ; tukan2023provable improves the training efficiency by adaptively selecting the best subset of training data (see App. A).
Concept-based interpretability approaches. Concepts are defined as high-level semantics that refers to the abstract and human-interpretable meanings of the visual data, such as objects, actions, and scenes, as opposed to low-level features like edges or textures wu2016valueexplicithighlevel . Concepts have been used in interpretable computer vision to bridge the gap between human understanding and machine perception in various tasks such as image classification. Such interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods (see App. A). Concept Bottleneck Models (CBMs) extend interpretable-by-design approaches by using human-understandable attributes as an intermediate layer for predictions, as used in few-shot learning lampert2013attribute and attribute learning xu2020attribute ; russakovsky2012attribute . While interpretable, CBMs reliance on costly annotations and lower accuracy compared to end-to-end models limit their usage. Post-hoc Concept Bottleneck Models (PCBMs) address these issues by incorporating static knowledge bases (e.g., ConceptNet speer2017conceptnet ) and residual connections to boost accuracy yuksekgonul2022post . Recently yang2023language ; yan2023learning incorporated LLMs to identify the concept bottleneck making classification more explainable. We build on this and use CBMs for CS.
3 Preliminaries
3.1 Coreset selection (CS) problem formulation
Consider a classification task and data distribution $P$. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the dataset of $n$ training examples sampled i.i.d. from the distribution $P$ where $x_{i}$ denotes the data and $y_{i}\in\mathcal{Y}$ denotes the label from a set of $N$ classes. CS coleman2019selection ; zheng2022coverage aims to find a subset $\mathcal{S}$ of $\mathcal{D}$ consisting of $m\leq n$ samples such that the models trained on $\mathcal{S}$ achieve performance comparable to models trained on $\mathcal{D}$. Let $\theta_{\mathcal{D}}$ and $\theta_{\mathcal{S}}$ denote the “downstream model" trained on $\mathcal{D}$ and $\mathcal{S}$ (coreset), respectively and $\ell$ be the loss function then the CS problem is as follows
$\min_{\mathcal{S}:\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=m}\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{S}})]-\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{D}})].$ (1)
To find this subset $\mathcal{S}$, previous works have proposed scores that gauge a sample’s difficulty for a model. Approaches such as max entropy uncertainty sampling lewis1994heterogeneous ; Settles_2012 , and least confidence culotta2005reducing estimate difficulty using the uncertainty of the model’s predictions on a sample. Another set of approaches such as $k$-center greedy sener2017active uses geometric information of the data to filter out redundant samples. Yet, another set of approaches uses training dynamics of the downstream model to estimate the difficulty score. Scores such as the forgetting events toneva2018empirical computed as the number of times a sample gets misclassified after being correctly classified earlier during model training, and the area under the margin (AUM) pleiss2020identifying which identifies mislabeled/difficult samples, fall in this category.
3.2 Concept bottleneck models (CBMs)
Recent advances in language model-guided CBMs utilize an off-the-shelf LLM to obtain concept bottlenecks which are then used to predict the labels. These works rely on pre-trained multi-modal models (such as CLIP radford2021learning ) which consists of a visual encoder $\mathcal{V}_{enc}$ and a text encoder $\mathcal{T}_{enc}$ that can map images and text to a $d$-dimensional representation space. Let $C=\{c_{1},c_{2},\cdots,c_{N_{C}}\}$ be the set of $N_{C}$ concepts (bottleneck) generated via an LLM, we can then construct a bottleneck embedding matrix $E_{C}\in\mathcal{R}^{N_{C}\times d}$ such that each row of the matrix is mapping of the concept $c\in C$ after passing it through textual encoder $\mathcal{T}_{enc}$. Based on this, a CBM yang2023language produces a prediction $h(x)=f(g(\mathcal{V}_{enc}(x);E_{C}))$ for a sample $x$ where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N_{C}}$ computes the similarity of the visual features to each concept in the bottleneck and $f:\mathbb{R}^{N_{C}}\rightarrow\Delta$ outputs the probability of each class in the label set $\mathcal{Y}$, where $\Delta$ is a $N$ simplex. We discuss details of $f$ and $g$ in Sec. 4.
4 Methodology
4.1 Generating the concept bottleneck via off-the-shelf LLMs
Since obtaining data with concept annotation is costly, we use LLMs to generate concept annotation for the samples. However, generating attributes (word-level concepts) for all the samples in the dataset via LLMs is still costly, hence we generate word-level concepts only for class label names. This approach was recently shown to be effective at generating the concept bottleneck for interpretable image classification  yan2023learning ; yang2023language . In Figure 2 (block 1), we present the prompts provided to the LLMs to extract the concepts for various class label names. The responses of the LLM are then processed to remove formatting errors to obtain the concept sets. Once the per-class concepts are extracted, we select $k$ discriminative concepts per class (concepts that are unique to a class) to form the concept bottleneck. Our final list of $k$ concepts for a class contains its class name and $k-1$ attributes generated by the LLM. Details of our prompt design, robustness check of different prompts, and examples of the generated attributes are mentioned in App. B.
4.2 Concept-based score for CS
We now describe how to use the concept bottleneck to produce a difficulty score for the samples in the dataset. We start by discussing how we learn the functions $f$ and $g$ described in Sec. 3.2 (see Fig. 2 (block 2)). We use the dot product between the visual embeddings of an image $x$ i.e. $\mathcal{V}_{enc}(x)$ and the bottleneck embedding matrix $E_{C}$ to measure the alignment between the visual and textual features yang2023language ; yan2023learning . Concretely, we compute the concept similarity score for a sample $x$ as
$g(x;E_{C}):=\mathcal{V}_{enc}(x)\cdot E_{C}^{\intercal}.$ (2)
To map the concept similarity score to a prediction in the label space $\mathcal{Y}$, we propose to use a linear (concept bottleneck layer) predictor denoted by $f$. Concretely, the function $f$ with parameters $W\in\mathbb{R}^{N\times N_{C}}$ is given by $f(x;W):=g(x;E_{C})\cdot W^{\intercal}$. We learn the parameters $W$ using
$W^{*}=\arg\min_{W}\frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i};W),y_{i}),$ (3)
where $\ell(f(x;W),y)=-\log(f(x;W)_{y})$ is the cross-entropy loss. The output of the concept bottleneck layer is defined as, $h(x):=f(g(x;E_{C});W^{*}).$ In practice, we learn $W$ using mini-batch gradient descent by running the optimization for $T$ epochs. We then compute the difficulty of each training sample using area under the margin (AUM) pleiss2020identifying while solving Eq. 3, quantifying the data difficulty as the margin of a training sample averaged over $T$ training epochs. Concretely, margin of a sample $(x,y)$ at an epoch $t$ is $M^{t}(x,y)=h^{t}_{y}(x)-\max_{y^{\prime}\neq y}h^{t}_{y^{\prime}}(x),$ where $h^{t}_{y^{\prime}}(x)$ is the prediction likelihood of the bottleneck layer at epoch $t$ for a class $y^{\prime}$. Thus, AUM (concept-based score) is
$\mathrm{AUM}(x,y)=\frac{1}{T}\sum_{t=1}^{T}M^{t}(x,y).$ (4)
Recent works pleiss2020identifying ; zheng2022coverage ; zheng2024elfs have demonstrated the effectiveness of AUM for computing a sample’s difficulty for CS. However, zheng2022coverage ; zheng2024elfs compute AUM for a specific downstream model by training it on the entire dataset first, which is computationally costly. On the other hand, we integrate AUM with the training of a linear layer $h$, training which is significantly cheaper than training the downstream model (training $h$ for 100 epochs takes only $7$ minutes on Imagenet compared to $8$ hours for training a ResNet-34 for 100 epochs). Moreover, since our score is independent of the downstream model, our coresets can be used for any downstream model without change, unlike training dynamics-based approaches that require computing their coresets again for different/new downstream models.
Sampling training examples to form a coreset. After obtaining data difficulty scores, a crucial step is choosing the samples to form the coreset. While many previous works toneva2018empirical ; coleman2019selection have reported encouraging results keeping the most challenging samples (for our concept-based score this means samples with the smallest margin), recent works zheng2022coverage ; sorscher2022beyond have shown that this could lead to a catastrophic drop in accuracies after training the downstream model on the coreset, especially when the size of the coreset is small. This is mainly due to poor sample coverage and potentially mislabeled data in the datasets. To remedy this, we use Coverage-centric Coreset Selection (CCS) proposed by zheng2022coverage (see Alg. 1 in App. C.5) which filters out (potentially) mislabeled samples and uses a stratified sampling approach to form the coreset. This technique has been shown to consistently achieve superior results to the random baselines for various coreset sizes.
4.3 Concept-based score for label-free CS
Recently, there has been an interest zheng2024elfs ; maharana2023d2 ; griffin2024zero in identifying the representative samples from an unlabeled dataset so as to 1) reduce the samples that need to be labeled/annotated by humans and 2) improve the efficiency of model training by only training the model on a subset of data. Our concept-based score can also be effectively utilized for this task with a simple modification. Similar to previous works maharana2023d2 ; zheng2024elfs ; sorscher2022beyond , we assume that we know the number of classes in the datasets. Additionally, we assume that we also know the names of the classes in the datasets. Previous works have demonstrated that VLMs such as CLIP radford2021learning achieve excellent zero-shot performance without requiring fine-tuning on specific datasets. We leverage this capability of CLIP models to obtain pseudo-labels for our unlabeled dataset and use them to obtain our difficulty score as follows
$\mathrm{AUM}(x,y_{\texttt{pseudo}})=\frac{1}{T}\sum_{t=1}^{T}M^{t}(x,y_{\texttt{pseudo}}),$ (5)
where for an image $x$ in the dataset $y_{\texttt{pseudo}}=\arg\max_{j\in\mathcal{Y}}\mathcal{V}_{enc}(x)\cdot\mathcal{W}^{\intercal}_{\texttt{zeroshot}}$ where $\mathcal{W}_{\texttt{zeroshot}}\in\mathbb{R}^{N\times d}$ is a matrix with columns defined as $\mathcal{T}_{enc}(s_{j})$ and $s_{j}=$ “a photo of a {$j^{th}$ class name}" for each class $j\in\mathcal{Y}$ radford2021learning ; wortsman2022robust . We use these scores along with CCS zheng2022coverage to produce the coreset. Similar to zheng2024elfs ; maharana2023d2 , the coreset is then assumed to be annotated by humans and models are trained on it.
5 Experiments
Datasets, models, and training: We focus on CS for classification tasks on three benchmark datasets namely, CIFAR-10, CIFAR-100 krizhevsky2009learning , and Imagenet-1K deng2009imagenet datasets consisting of $50000$, $50000$, and $1.28$ million samples spread across $10$, $100$, and $1000$ classes, respectively. For CIFAR-10/CIFAR-100, we train a ResNet(RN)-18 as a downstream model and for Imagenet we train ResNet-18/34/50 he2016deep , MobileNet sandler2018mobilenetv2 , DenseNet huang2017densely , Wide ResNet zagoruyko2016wide , and ViT dosovitskiy2020image as downstream models on the coresets. We use FFCV leclerc2023ffcv to accelerate training on Imagenet. We run CS for three trials with different random seeds and report the average of these runs in our tables for various pruning rates where a pruning rate of $90$% refers to removing $90$% of the samples.
For generating concepts we use a recently proposed open source model LLaVA liu2023llava ; liu2023improvedllava . For computing the concept similarity scores between the visual and concept features we used the CLIP radford2021learning model (with the ViT B-32 as backbone) following the previous works yun2022vision ; yang2023language ; yan2023learning . For computing the pseudo-labels for label-free CS (in Sec. 4.3) we used a ViT L-14 CLIP model trained on the DataComp-1B dataset ilharco_gabriel_2021_5143773 . Further experimental details are mentioned in App. C.
Table 1: Comparison of the model’s (RN-18 for CIFAR10/100 and RN-34 for Imagenet) test accuracy after training on coresets found by various approaches shows that our coresets lead to significantly better performance than Random achieving competitive results compared to the methods using the downstream model’s training dynamics, even for high pruning rates. (Best in each category is highlighted).
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100 Imagenet
$30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$
Needs downstream model training/dynamics Entropy $94.44$ $92.11$ $85.67$ $66.52$ $72.26$ $63.26$ $50.49$ $28.96$ $72.34$ $70.76$ $64.04$ $39.04$
Forgetting $\mathbf{95.40}$ $\mathbf{95.04}$ $92.97$ $85.70$ $\mathbf{77.14}$ $\mathbf{74.45}$ $\mathbf{68.92}$ $\mathbf{55.59}$ $\mathbf{72.60}$ $\mathbf{70.89}$ $66.51$ $52.28$
AUM $95.27$ $94.93$ $\mathbf{93.00}$ $\mathbf{86.08}$ $76.84$ $73.77$ $68.85$ $55.03$ $72.29$ $70.52$ $\mathbf{67.78}$ $\mathbf{57.36}$
Does not need downstream model training/dynamics Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
RandomFFCV - - - - - - - - $73.37$$\pm 0.08$ $71.71$$\pm 0.10$ $67.85$$\pm 0.04$ $51.29$$\pm 0.20$
Ours $\mathbf{94.77}$$\pm 0.09$ $\mathbf{93.44}$$\pm 0.61$ $\mathbf{91.80}$$\pm 0.21$ $\mathbf{84.63}$$\pm 0.24$ $\mathbf{75.98}$$\pm 0.26$ $\mathbf{72.22}$$\pm 0.22$ $\mathbf{66.53}$$\pm 0.42$ $\mathbf{51.85}$$\pm 0.29$ $\mathbf{73.39}$$\pm 0.12$ $\mathbf{72.34}$$\pm 0.13$ $\mathbf{69.44}$$\pm 0.17$ $\mathbf{55.92}$$\pm 0.02$
Table 2: Comparison of the model’s test accuracy after training on coresets, found in a label-free manner, shows that our coresets lead to better performance than Random and Prototypicality. It is also competitive/better than the coresets found by ELFS (which uses training dynamics of the downstream model).
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100 Imagenet
$30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$ $30\%$ $50\%$ $70\%$ $90\%$
Needs downstream model training/dynamics ELFS (SwAV) $95.00$ $94.30$ $91.80$ $82.50$ $76.10$ $72.10$ $65.50$ $49.80$ $73.20$ $71.40$ $66.80$ $53.40$
ELFS (DINO) $\mathbf{95.50}$ $\mathbf{95.20}$ $\mathbf{93.20}$ $\mathbf{87.30}$ $\mathbf{76.80}$ $\mathbf{73.60}$ $\mathbf{68.40}$ $\mathbf{54.90}$ $\mathbf{73.50}$ $\mathbf{71.80}$ $\mathbf{67.20}$ $\mathbf{54.90}$
Does not need downstream model training/dynamics Prototypicality $94.70$ $92.90$ $90.10$ $70.90$ $74.50$ $69.80$ $61.10$ $32.10$ $70.90$ $60.80$ $54.60$ $30.60$
Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
RandomFFCV - - - - - - - - $73.37$$\pm 0.08$ $71.71$$\pm 0.10$ $67.85$$\pm 0.04$ $51.29$$\pm 0.20$
Ours-LF $\mathbf{94.81}$$\pm 0.14$ $\mathbf{93.93}$$\pm 0.13$ $\mathbf{91.75}$$\pm 0.34$ $\mathbf{84.02}$$\pm 0.44$ $\mathbf{74.67}$$\pm 0.23$ $\mathbf{72.07}$$\pm 0.58$ $\mathbf{65.50}$$\pm 0.17$ $\mathbf{49.91}$$\pm 0.96$ $\mathbf{73.61}$$\pm 0.08$ $\mathbf{71.99}$$\pm 0.05$ $\mathbf{68.42}$$\pm 0.21$ $\mathbf{53.21}$$\pm 0.06$
CS baselines and methods: We compare our method against various baselines and SOTA CS methods. 1) Random: Uniformly select samples from the datasets to form the coreset. RandomFFCV denotes the performance of the models trained on random subsets of Imagenet using FFCV leclerc2023ffcv . 2) Entropy coleman2019selection : Selects samples based on entropy computed as the uncertainty of a model’s prediction on a sample. 3) Forgetting toneva2018empirical : Selects samples based on the forgetting score computed as the number of times a sample is misclassified after being correctly classified earlier during training of the downstream model. A higher forgetting score indicates a more challenging sample. 4) AUM pleiss2020identifying : Selects samples based on their average margin during training of the downstream model i.e., the difference between the target class and the next highest class across the training epochs. Lower AUM indicates a more challenging sample. For forgetting, AUM, and our method, we use CCS zheng2022coverage to form the coreset whereas for entropy we select samples with the highest entropy as done in coleman2019selection ; zheng2022coverage .
For label-free CS, we use 1) Prototypicality sorscher2022beyond : which first performs k-means clustering in the embedding space of SwAV caron2020unsupervised model and ranks samples based on their Euclidean distance to the cluster centers. Samples further away from the cluster center are used to form the coreset. 2) ELFS zheng2024elfs : estimates the pseudo-labels of the unlabeled samples using a deep clustering approach (using the embedding space of SwAV caron2020unsupervised and DINO caron2021emerging ) and forms the coreset using training dynamics of the downstream model trained on the pseudo-labeled data. Crucially, SOTA methods such as forgetting, AUM, and ELFS require training the downstream model on the entire dataset first for CS, unlike our method which is independent of the downstream model. While Random and Prototypicality also don’t require the downstream model for CS, our results show that our approach surpasses them.
5.1 Evaluating performance of our concept-based score for CS
Table 1 shows the accuracy of models trained on coresets found by various approaches on three datasets for the standard CS problem (where the dataset is labeled). Our results show that our coresets lead to significantly better performance, even at higher pruning rates, compared to the random subsets. Moreover, our method which does not have the knowledge of the downstream model or its training dynamics still provides competitive performance to coresets found by the SOTA approaches based on forgetting and AUM, and even outperforms them on Imagenet for smaller pruning rates.
Table 2 shows the accuracy of models trained on coresets when the training set is unlabeled (we report the numbers presented by zheng2024elfs for previous methods). The results show that the random subsets are a competitive baseline for this problem outperforming Prototypicality sorscher2022beyond . Our results also show that our coresets outperform the random subsets for all pruning rates with significant improvements at higher pruning rates. Compared to ELFS zheng2024elfs , our method provides competitive performance and even surpasses it for lower pruning rates on Imagenet, without using any information about the downstream model’s architecture or its training dynamics highlighting its effectiveness.
Next, evaluate the transferability of the coresets found by our approach to downstream models with different architectures including those based on convolutional neural networks, ResNets and ViTs. Results in Table 3 and Table 13 (in App. C.4) show that our coresets achieve superior performance than random for all the architectures. Thus, our concept-based score in conjunction with stratified sampling zheng2022coverage which balances the number of easy/hard examples in the coreset is an effective approach for both standard and label-free CS at various pruning rates.
Figure 3: Visualizing samples according to our concept-based score for a subset of classes in CIFAR-10/100 showing that easy (challenging) samples are aligned (unaligned) with their assigned label. Image-level concepts (in boxes) extracted via LLaVA confirm that easy (challenging) examples are aligned (unaligned) with concepts of their labels, explaining the reason for a high (low) concept-based score.
5.2 Evaluating efficiency of CS via our concept-based approach
Here we compare the efficiency of our approach at finding coresets compared to approaches based on training dynamics. For training dynamics-based approaches the time required to find the coreset is dominated by the time needed to train the downstream model on the entire dataset first. For example, finding a coreset of Imagenet using a ResNet-34 model takes roughly 8 hours on two A-100 GPUs. In comparison, for our approach, extracting concepts via LLMs (block 1 of Fig. 2) takes 3 seconds per class (totaling to 25 minutes for all classes of Imagenet using 2 GPUs), and computing visual/concept features via CLIP and training a linear concept bottleneck layer (block 2 for Fig. 2) takes roughly 30 minutes for Imagenet offering $\approx 8$x speedup for CS over approaches relying on training dynamics of the downstream model. Moreover, since our method is independent of the downstream model architecture we do not need to repeat the CS step for different architectures. This is in contrast with training dynamics-based methods that require training the downstream model for every new architecture to find the corresponding best coreset for it.
Table 3: Superior performance of downstream models with various architectures trained on our coresets for Imagenet compared to Random for standard (Ours) and label-free (Ours-LF) CS highlight our method’s effectiveness for CS without knowledge of the downstream model.
Model Architecture Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
RN-18 Random $71.15_{\scriptscriptstyle\pm 0.23}$ $68.48_{\scriptscriptstyle\pm 0.10}$ $63.15_{\scriptscriptstyle\pm 0.19}$ $44.96_{\scriptscriptstyle\pm 0.50}$
Ours-LF $71.21_{\scriptscriptstyle\pm 0.09}$ $68.77_{\scriptscriptstyle\pm 0.13}$ $63.76_{\scriptscriptstyle\pm 0.09}$ $47.50_{\scriptscriptstyle\pm 0.10}$
Ours $70.94_{\scriptscriptstyle\pm 0.19}$ $69.30_{\scriptscriptstyle\pm 0.84}$ $65.16_{\scriptscriptstyle\pm 0.04}$ $49.57_{\scriptscriptstyle\pm 0.16}$
RN-50 Random $76.06_{\scriptscriptstyle\pm 0.11}$ $74.44_{\scriptscriptstyle\pm 0.04}$ $70.50_{\scriptscriptstyle\pm 0.02}$ $53.56_{\scriptscriptstyle\pm 0.13}$
Ours-LF $76.54_{\scriptscriptstyle\pm 0.08}$ $74.84_{\scriptscriptstyle\pm 0.03}$ $71.10_{\scriptscriptstyle\pm 0.09}$ $55.22_{\scriptscriptstyle\pm 0.92}$
Ours $76.29_{\scriptscriptstyle\pm 0.10}$ $75.12_{\scriptscriptstyle\pm 0.03}$ $72.05_{\scriptscriptstyle\pm 0.09}$ $58.26_{\scriptscriptstyle\pm 0.46}$
DenseNet Random $76.55_{\scriptscriptstyle\pm 0.11}$ $75.31_{\scriptscriptstyle\pm 0.15}$ $71.85_{\scriptscriptstyle\pm 0.20}$ $56.27_{\scriptscriptstyle\pm 0.21}$
Ours-LF $76.94_{\scriptscriptstyle\pm 0.23}$ $75.52_{\scriptscriptstyle\pm 0.11}$ $72.28_{\scriptscriptstyle\pm 0.14}$ $58.16_{\scriptscriptstyle\pm 0.04}$
Ours $76.80_{\scriptscriptstyle\pm 0.22}$ $75.94_{\scriptscriptstyle\pm 0.03}$ $73.40_{\scriptscriptstyle\pm 0.27}$ $60.93_{\scriptscriptstyle\pm 0.19}$
Wide Resnet Random $77.39_{\scriptscriptstyle\pm 0.19}$ $75.32_{\scriptscriptstyle\pm 0.12}$ $71.51_{\scriptscriptstyle\pm 0.10}$ $54.78_{\scriptscriptstyle\pm 0.83}$
Ours-LF $77.83_{\scriptscriptstyle\pm 0.08}$ $75.77_{\scriptscriptstyle\pm 0.09}$ $71.97_{\scriptscriptstyle\pm 0.08}$ $57.34_{\scriptscriptstyle\pm 0.58}$
Ours $77.48_{\scriptscriptstyle\pm 0.05}$ $76.32_{\scriptscriptstyle\pm 0.03}$ $73.14_{\scriptscriptstyle\pm 0.03}$ $59.75_{\scriptscriptstyle\pm 0.70}$
5.3 Visualizing easy and challenging samples based on our concept-based score
Here we show the advantage of our concept-based score for assessing the sample’s difficulty and how using concepts aids its interpretability. We start by visualizing the easiest and the most challenging images (per class) for CIFAR-10/100. In Fig. 3, the top row shows the images with the highest concept-based scores (easiest) and the bottom shows the images with the lowest scores (challenging) for a subset of classes in CIFAR-10/100. As observed the easiest images are typical images associated with the label where as the challenging images are confusing (and even potentially mislabeled) as they look like images from a different class. For example, some challenging images in the class “boy” from CIFAR-100 are actually images of a baby which is also a class in CIFAR-100. Similarly, some challenging images from the class “cat” in CIFAR-10 look like images of a dog. More examples of such images are presented in Fig. 4 in App. C.1. Since the challenging examples seem to be confusing, it shows that our score can identify examples that are ambiguous or mislabeled. Such samples may be hard for some ML models to learn and could force them to rely on spurious features, potentially hurting generalization. The ability of our score to identify such samples without the downstream model demonstrates its effectiveness for ranking the samples for CS.
Next, we demonstrate why certain samples get low/high concept-based scores in our approach by extracting concepts specific to these images using LLaVA (note that these concepts are different from the per-class concepts used in the concept bottleneck). To generate these, we prompt LLaVA to produce concepts using both the sample image and its class label (see image-level concept extraction in App. B). These image-level concepts are shown in the boxes in Fig. 3. As observed in Fig. 3, image-level concepts provided by LLaVA are related to the class label for easy images whereas they are unrelated for challenging images. For example, attributes provided by LLaVA for the challenging images of “airplane” align more with those of ship (both which are classes in CIFAR-10), and concepts provided for challenging images of “bridges” align more with those of castles (both of which are classes in CIFAR-100). Since our concept-based score in Eq. 4 assigns a small value for these images, we see that it correctly captures when a sample’s visual information is not aligned with the sample’s associated label and vice-versa. Thus, explaining why certain examples should be included/excluded from the coreset in an interpretable way independent of the downstream model.
5.4 Ablation studies
Here, we present ablation studies to evaluate various component of our approach. Specifically, we evaluate the effect of different 1) methods and LLMs used for concept generation, 2) number of concepts per class ($k$) in the bottleneck, 3) CLIP backbones for visual/concept similarity, 4) number of iterations ($T$) for training the concept bottleneck layer. Here we use CIFAR-100 with a $90\%$ pruning ratio and train a ResNet-18 on the selected coresets. Lastly, we also evaluated the effectiveness of our method for CS on tasks beyond object recognition using medical and affective computing tasks.
Comparison of different techniques to generate concepts via LLaVA. Tables 2 and 5 shows how the performance of models trained on our coresets change when different method are used to generate the concept sets (Fig. 2 block 1). Since our method uses LLaVA liu2023llava , which is a VLM, we compare the performance of models trained on the random subsets and coresets obtained using class-wise concepts (only textual information) and concepts extracted using both visual and textual information.
Table 4: Effect of using concepts obtained using only textual information (attributes vs. descriptions) to form the concept bottleneck layer  [^2] Cf. A random subset of CIFAR-100 achieves an accuracy of $44.76_{\pm 1.58}$ at 90% pruning rate..
class-wise attributes class-wise descriptions
Ours $51.85_{\scriptscriptstyle\pm 0.51}$ $51.05_{\scriptscriptstyle\pm 0.71}$
Ours-LF $49.91_{\scriptscriptstyle\pm 0.96}$ $49.85_{\scriptscriptstyle\pm 0.83}$
For concepts generated using only textual information, we consider two alternatives, namely class-wise attributes (CW-A) and class-wise descriptions (CW-D). While CW-A considers concepts formed by a single or a few words, CW-D consists of longer, more descriptive concepts (eg., a descriptive concept for the class butterfly is “a beautiful insect with colorful wings"). For CW-D, we use a subset of $k$ concepts provided by yang2023language , generated via the GPT-3 model. Our results show that CW-A performs better than CW-D for both the standard/label-free CS problems. Thus, we used CW-A for all our experiments.
Table 5: Effect of using concepts generated by using both visual and textual information to form the concept bottleneck layer footnotemark: .
class-wise one-shot image attributes image-wise attributes
Ours $51.68_{\scriptscriptstyle\pm 0.45}$ $52.47_{\scriptscriptstyle\pm 0.39}$
Ours-LF $50.22_{\scriptscriptstyle\pm 0.16}$ $51.05_{\scriptscriptstyle\pm 0.93}$
Next, for generating concepts using both visual and textual information, we consider two alternatives. The first is a class-wise one-shot image attribute approach where we first cluster all images of a class in the embedding space of the CLIP’s visual encoder and identify the image whose embedding is the closest to the cluster center (for the label-free setting we use the pseudo-labels of the images during clustering), then we prompt LLaVA to generate attributes using this single image and the class name. Once generated, we use $k$ discriminative concepts to form the bottleneck. The second alternative is the image-wise attribute approach, where we use each image in the training set and prompt LLaVA to generate per image attributes describing the image. Once generated, we sort the concepts based on their frequency of occurrence in a class and use the most frequently occurring discriminative concepts to form the bottleneck. While the image-level concepts lead to the best coresets, it is slow and costly to prompt LLaVA to generate attributes for all the images in a large dataset such as Imagenet.
Table 6: Effect of the number of per-class concepts $k$ and the method of selecting $k$ attribute-level concepts from those generated by LLaVA on the accuracy of models trained on the coresetfootnotemark: .
Concept selection $k$=1 $k$=5 $k$=10
Random concepts $48.78_{\scriptscriptstyle\pm 0.96}$ $50.15_{\scriptscriptstyle\pm 1.64}$ $50.39_{\scriptscriptstyle\pm 0.72}$
Discriminative $51.42_{\scriptscriptstyle\pm 0.18}$ $\mathbf{51.85}_{\scriptscriptstyle\pm 0.29}$ $51.22_{\scriptscriptstyle\pm 0.72}$
For CIFAR-100, this process took about nine hours to complete (in comparison CW-A can be extracted in 5 minutes without parallel computation) which is very costly compared to the small performance gains it provides over other approaches. Lastly, while the class-wise one-shot image attribute approach is better than CW-A, the additional step of clustering can be costly for larger datasets such as Imagenet. Based on these results, we used CW-A for concept generation using LLaVA.
Effect of $k$ and the method for selecting $k$ concepts. In Table 6, we show how the number of concepts extracted per class label, for creating the bottleneck in block 1 of Fig. 2 affects the selection of coresets. Once the list of class-wise attribute-level concepts is generated by LLaVA, we can select $k$ concepts per class either randomly or choose concepts unique to a class (discriminative). Our results show that using even $k=1$ is sufficient to surpass the performance using a random subsetfootnotemark: . This performance increases when we keep discriminative concepts in our concept bottleneck, with $k=5$ achieving the best results. While the size of the concept bottleneck need not be large to find good coresets, it is helpful to take a sample’s visual similarity with a set of concepts rather than a single concept per class. Thus, we used $5$ discriminative concepts per class to form the bottleneck.
Table 7: Effect of using different LLMs for generating concepts used in the concept bottleneckfootnotemark: .
LLM Accuracy
GPT-3 $51.05_{\scriptscriptstyle\pm 0.71}$
Phi-3 $51.30_{\scriptscriptstyle\pm 0.26}$
LLaVA $51.85_{\scriptscriptstyle\pm 0.29}$
Table 8: Effectfootnotemark: of using various CLIP backbones for computing the visual-concept similarity in Eq. 2.
CLIP backbone Accuracy
ViT B-16 $48.89_{\scriptscriptstyle\pm 1.01}$
ViT L-14 $49.54_{\scriptscriptstyle\pm 1.82}$
ResNet-50 $50.56_{\scriptscriptstyle\pm 1.67}$
ViT B-32 $51.85_{\scriptscriptstyle\pm 0.29}$
Table 9: Effectfootnotemark: of using different number training epochs $T$ for computing the score in Eq. 4.
$T$ Accuracy
10 $46.63_{\scriptscriptstyle\pm 0.08}$
20 $48.40_{\scriptscriptstyle\pm 0.10}$
50 $50.50_{\scriptscriptstyle\pm 0.31}$
100 $51.85_{\scriptscriptstyle\pm 0.29}$
Effect of using different LLMs/VLMs for concept extraction. We evaluated three different VLMs for concept extraction in block 1 of Fig. 2. We used GPT-3 and two open source VLMs namely Phi-3-Mini-4K-Instruct with 3.8 billion parameters, and LLaVA with 7 billion parameters. In Table 9, we find that the performance of our method remains stable regardless of the LLM used, indicating that even smaller LLMs are effective at generating concepts that produce high performing coresets.
Effect of using different VLMs for measuring visual and concept similarity. Here we evaluated different CLIP backbones to compute the similarity between the visual and concept features used in Eq. 2 (block 2 of Fig. 2). Our results in Table 9 show that our concept-based method achieves significantly better performance than Randomfootnotemark: for all the backbones with ViT B-32 backbone performing the best. Thus, we used this backbone for all the experiments in the paper.
Effect of number of epochs $T$. Here we evaluated the effect of using different number of epochs $T$ used for training the concept bottleneck layer for computing the concept-based score. Our results in Table 9 show that $T\geq 50$ is enough to achieve concept-based scores that lead to selection of high performing coresets. We used $T=100$ in our experiments since that achieves the best performance.
Evaluation on medical and affective computing tasks. In App. C.2, we present an evaluation of using our approach for CS for biomedical entity recognition (using BloodMNIST acevedo2020dataset ) and emotion recognition (using a subset of AffectNet mollahosseini2017affectnet ) task. Superior performance of our approach in Tables 10 and 11 on these tasks highlights the effectiveness of our approach for CS on diverse tasks.
6 Conclusion
CS finds representative samples from a large dataset, training models on which leads to models with accuracy similar to the models trained on the entire dataset. In this work, we proposed a scoring mechanism based on concept bottlenecks that allows us to compute the difficulty of a sample in terms of interpretable concepts. This method is independent of the downstream model and avoids training the downstream model on the full dataset even once. Our experiments show that training downstream models on coresets selected using our approach leads to better performance than random subsets and achieves accuracy similar to or better than the SOTA approaches based on training dynamics of the downstream model, for both the standard and label-free CS problem. Moreover, our score provides an intuitive explanation of a sample’s difficulty at a dataset level, independent of the downstream model.
Limitations: While we evaluated various LLMs for concept generation and showed their effectiveness for computing concept-based score, it is possible that for some classes LLMs produce very noisy and non-discriminative concepts, leading to poor concept-based scores. Thus, developing methods to prompt the LLMs to generate discriminative concepts to be used as the concept bottleneck is a crucial research direction and is left for future works. Another limitation is the dependence of our score on the VLMs for computing similarity between visual and concept features. Although we evaluated our approach on a diverse set of tasks including a biomedical task and various CLIP-based VLMs, it is possible that these similarity scores may not work well for certain tasks, warranting additional fine-tuning to make the similarity scores meaningful. Hence, in practice, additional time/compute may be required to identify or fine-tune a VLM for CS.
Appendix
We present additional related work in Appendix A. Then we describe the details of our methodology for extracting the concepts from LLaVA in Appendix B and present additional experiments and implementation details of our experiments in Appendix C including the algorithm for stratified sampling used in our work in Appendix C.5.
Appendix A Additional related work
Concept-based interpretability: Interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Post-hoc methods include Gradient-weighted Class Activation Mapping approaches [3, 59, 46, 23] that trace network gradients to identify the input areas that guide predictions and Explanation Generation methods [63, 48, 32, 20] that require models to produce explanations for visual tasks by conditioning their predictions on captioning models or incorporating visual evidence to ground explanations [21, 49]. Interpretable-by-design methods, such as Prototype methods, optimize a metric space where classifications are based on distances to class prototypes, identifying important input regions but often obscuring their semantic content [47, 9, 64, 57, 72].
Concept Bottleneck Models (CBMs) are a part of interpretable-by-design approaches that use human-understandable attributes as an intermediate layer for predictions. A recent advancement, Computational Derivation Learning (CompDL), utilizes a CBM architecture by applying a linear layer over CLIP scores between expert-designed concepts and images, improving evaluation of how well CLIP grounds concepts [84]. Post-hoc Concept Bottleneck Models (PCBMs) were recently proposed to ease the requirement of CBMs to rely on costly concept annotations and improve their accuracy compared to end-to-end models. However, PCBMs are limited by the coverage of knowledge bases, making them unsuitable for large-scale or domain-specific tasks or fine-grained classification, and their residual predictors can undermine interpretability by blending CBMs with end-to-end models.
High-level semantic-driven descriptions are also used to guide data augmentation to build an informative set [74] to make model training efficient with a good enough training set. Prior works use external knowledge bases to obtain these textual semantic concepts to guide vision models [5, 29, 54, 62]. Thus, the use of concepts has been shown to improve interpretability in various domains. However, to the best of our knowledge, we are the first ones to propose a concept-based score for the CS problem and show its competitiveness to SOTA model training dynamics-dependent approaches.
Adaptive subset selection: These works focus of improving the training convergence and efficiency of model training by selecting a new subset of data from the whole dataset, every epoch, while training a downstream model. Thus, unlike our work, these works do not prune the dataset but rather keep selecting small, potentially non-overlapping, subsets every few epochs. Both [31, 71], propose ways to select subsets every epoch or every few epochs in a way that does not use a downstream model unlike some other works such as GradMatch [30], which select subsets based on the downstream model. While our work does not target adaptive subset selection, we evaluated how well our approach performs without any change on this problem in App. C.3. Our results show that, even on for this application, our concept-based score is an effective method.
Appendix B Details for concept set generation
[…]
Alternative VLMs for Concept Generation: We leverage LLaVA as our choice of VLM for concept generation, however in Table 9, we also compared against concepts extracted from GPT [79] and another smaller open source VLM (Phi-3-Mini-4K-Instruct). We see comparable performance of our CS method with concepts extracted from these models. We also experimented with retrieving concepts via another recent method, SpLiCE [4] which uses a linear optimization for sparse concept decomposition. However, a major limitation of SpLiCE is that similar to image-wise attributes as used in Table 5 it is a costly approach (SpLiCE can take up to 3 hours for $50,000$ images, significantly slower than generating class-level concepts from LLaVA).
Appendix C Additional experiments and implementation details
C.1 Visualizing easy/challenging samples based on concept-based score
Similar to Fig 3 in Sec. 5.3 of the main paper, we visualize easy and challenging examples in Fig. 4 for CIFAR-10 and subset of classes from CIFAR-100. As observed the easy images (the ones that get high scores in our approach) are more canonical images of the class labels whereas the challenging ones are images that can potentially be assigned another class in the same dataset or are mislabeled in the dataset. The clear distinctions between these images show that our concept-based score aligns well with human intuition on the difficulty of the samples.
Table 10: Performance of concept-based coreset selection for emotion recognition task on Affectnet dataset. Coresets selected by our approach results in superior performance (higher F1 score) compared to Random for standard CS.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
EfficientNet Random $0.607_{\scriptscriptstyle\pm 0.006}$ $0.573_{\scriptscriptstyle\pm 0.025}$ $0.517_{\scriptscriptstyle\pm 0.006}$ $0.347_{\scriptscriptstyle\pm 0.055}$
Ours $0.603_{\scriptscriptstyle\pm 0.006}$ $0.577_{\scriptscriptstyle\pm 0.021}$ $0.537_{\scriptscriptstyle\pm 0.015}$ $0.450_{\scriptscriptstyle\pm 0.035}$
Table 11: Performance of concept-based coreset selection for biomedical image recognition task on BloodMNIST dataset. Coresets selected by our approach results in superior accuracy compared to Random for standard CS, highlighting our method’s effectiveness across various tasks.
Model Arch. Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
ResNet-18 Random $94.99_{\scriptscriptstyle\pm 0.27}$ $94.79_{\scriptscriptstyle\pm 0.18}$ $91.71_{\scriptscriptstyle\pm 0.29}$ $86.50_{\scriptscriptstyle\pm 0.66}$
Ours $94.64_{\scriptscriptstyle\pm 0.60}$ $94.06_{\scriptscriptstyle\pm 0.10}$ $92.26_{\scriptscriptstyle\pm 0.19}$ $87.44_{\scriptscriptstyle\pm 1.46}$
C.2 Concept-based CS for emotion recognition and biomedical image recognition
To validate the effectiveness of our concept-based coreset selection method beyond object recognition tasks, we apply our concept-based CS approach to the task of emotion recognition and biomedical image recognition.
For emotion recognition, we use the Affectnet dataset [45] for our experiments. AffectNet is a large-scale facial expression dataset designed for training and evaluating affective computing models [73]. It contains facial images collected from the internet using web search queries for emotion-related keywords in multiple languages. Each image is manually annotated for eight discrete emotion categories: neutral, happiness, sadness, surprise, fear, disgust, anger, contempt. For our experiments, we utilize an openly available version of this dataset  [^3] https://www.kaggle.com/datasets/noamsegal/affectnet-training-data, containing roughly $16000$ training and $14000$ testing samples.
According to our approach we first use LLaVA to extract concepts for the 8 emotion classes, using the following prompt, “What are the facial features that distinguish emotion class name from other emotion types. Focus on changes in eyes, nose, lips, eyebrows, mouth. Give the output separated by commas in a line.”. We get $5-10$ distinctive facial feature concepts for every emotion, for instance for emotion class happy, we get the following concepts, “wide open eyes”, “sparking eyes”, “smiling lips”, “open mouth”, “raised eyebrows”, “flushed cheeks”, “teeth barred”. We finally select $k=5$ discriminative concepts from this list.
To test coreset performance, we use the EfficientNet model [67] and report F1 scores for our coresets in Table 10. When compared against randomly selected coresets for the various pruning ratios, coresets selected via our concept-based approach achieve better performance.
For biomedical image recognition, we use the BloodMNIST [1] dataset from the MedMNIST [80, 81] which comprises of images of normal blood cells, captured from individuals without infection, hematologic or oncologic disease and free of any pharmacologic treatment at the moment of blood collection. It consists of a total of $17,092$ images and is organized into 8 classes (basophil,eosinophil,erythroblast,immature granulocytes(myelocytes, metamyelocytes and promyelocytes),lymphocyte,monocyte,neutrophil,platelet).
For this dataset, we first extract concepts for the 8 blood cell types via GPT using the following prompt, “What are the features that can distinguish blood cell class name from rest of the blood cell types on their size, shape, nucleus appearance, and the presence of granules in their cytoplasm”. We obtain 10 concepts for every blood cell type, for instance, for platelets, we get the following concepts, “Smallest blood component”, “No nucleus”, “Granules present”, “Irregular shape”, “Cytoplasmic fragments”, “Variable granule distribution”, “Oval to round shape”, “Small dense granules”, “Lacks chromatin”, “Compact cytoplasmic body”.
To test the coreset performance, we use a ResNet-18 model and report accuracy of our coresets in Table 11. Similar to other results, our method achieves better performance than randomly selected coresets for higher pruning rates and is competitive at lower pruning ratios. This is attributes to the difficulty of calculating concept similarity is the representation space of the CLIP model which is potentially unaware of the terminology used in the medical domain. While replacing CLIP with a VLM that is trained on medical domain can boost the performance of our method, our results highlight that even with access to such a model our approach is able to find better coresets than random subsets.
Our results on these two tasks highlight the versatility of our approach for model-agnostic approach to coreset selection, which is able to find coreset without requiring training the the downstream models on the entire dataset even once.
Table 12: Adaptive coreset selection (Results for RBFNN and GradMatchPB are taken from tables 1 and 3 of Tukan et. al.)
Method Datasets and Pruning Rates
CIFAR-10 CIFAR-100
$70\%$ $90\%$ $70\%$ $90\%$
GradMatchPB [30] $91.89$ $90.01$ $72.57$ $60.39$
RBFNN [71] 94.4494.44{94.44}94.44 91.4091.40{91.40}91.40 $73.48$ $64.59$
Adaptive Random $93.64_{\scriptscriptstyle\pm 0.21}$ $90.49_{\scriptscriptstyle\pm 0.41}$ ${71.70}_{\scriptscriptstyle\pm 0.30}$ ${61.24}_{\scriptscriptstyle\pm 0.37}$
Ours (adaptive) $92.61_{\scriptscriptstyle\pm 0.08}$ $89.95_{\scriptscriptstyle\pm 0.26}$ ${71.74}_{\scriptscriptstyle\pm 0.33}$ $63.71_{\scriptscriptstyle\pm 0.48}$
C.3 Evaluation on adaptive subset selection
In this section, we present a comparison of using our concept-based method for the problem of adaptive subset selection [30, 31, 71] (See App. A for a description of the problem). To test our approach, we followed the experimental setup of [71], of training a ResNet-18 model on CIFAR-10/100 for $300$ epochs and changing the subset every $20$ epochs. Unlike works in the line of adaptive subset selection which focus on selecting subsets that do not overlap or train models on easy subsets first before moving to difficult ones, we simply used our CCS-based selection approach [86] to identify new subsets every $20$ epochs for this task. Due to the randomness in CCS (Alg. 1), samples being selected to form the coreset changes.
In our table 12, we compare our method to previous methods and with an adaptive random selection method (where a random subset of data is selected every $20$ epochs), which has been suggested by [31] as a strong baseline for this line of work. We observe that our method without any modifications achieves comparable performance to existing adaptive subset selection methods showing its effectiveness for this problem as well. Moreover, sampling new subsets via CCS (Alg. 1) is as efficient as sampling a random subset; thus our approach is also an efficient solution for this problem. While we believe that incorporating better subset sampling techniques as suggested by [31] may boost the performance of our method on this problem, CCS-based sampling is already quite effective.
C.4 Transferability of coresets
In this section, we present the performance of three additional downstream model architectures after training on the coresets found by our approach. Similar to the results in Table 3, our results in Table 13 show that coresets found by our approach in both standard and label-free setting achieve performance better than the random subsets on these three architectures as well. We note that the performance of the ViT model is worse than the performance of other model architectures with and without pruning (ViT-B-16 achieves an accuracy of $\approx 62\%$ compared to $\approx 78\%$ with ResNet-50 on full Imagenet dataset) due to the ViT models being data hungry in nature [14] and the fact that we used standard SGD-based training (similar to that used for training other models in the paper). We believe that using other training methods for ViTs as suggested in [69] could produce better performing ViT models.
Nonetheless, better performance than random subsets across a variety of downstream model architectures highlights the effectiveness of our approach at finding coresets without the knowledge of the architecture or training dynamics of the downstream models.
Table 13: Superior performance of downstream models with different architectures trained on our coresets for Imagenet compared to Random for both standard (Ours) and label-free (Ours-LF) CS highlight our transferability of our coresets.
Model Architecture Method Pruning Rates
$30\%$ $50\%$ $70\%$ $90\%$
MobileNet Random $62.68_{\scriptscriptstyle\pm 0.09}$ $62.49_{\scriptscriptstyle\pm 0.15}$ $61.53_{\scriptscriptstyle\pm 0.29}$ $53.80_{\scriptscriptstyle\pm 0.17}$
Ours-LF $61.60_{\scriptscriptstyle\pm 0.18}$ $61.97_{\scriptscriptstyle\pm 0.08}$ $61.73_{\scriptscriptstyle\pm 0.25}$ $54.39_{\scriptscriptstyle\pm 0.66}$
Ours $61.36_{\scriptscriptstyle\pm 0.15}$ $62.32_{\scriptscriptstyle\pm 0.22}$ $62.46_{\scriptscriptstyle\pm 0.26}$ $55.57_{\scriptscriptstyle\pm 0.13}$
RN-34 Random $73.37_{\scriptscriptstyle\pm 0.08}$ $71.71_{\scriptscriptstyle\pm 0.10}$ $67.85_{\scriptscriptstyle\pm 0.04}$ $51.29_{\scriptscriptstyle\pm 0.20}$
Ours-LF $73.61_{\scriptscriptstyle\pm 0.08}$ $71.99_{\scriptscriptstyle\pm 0.05}$ $68.42_{\scriptscriptstyle\pm 0.21}$ $53.52_{\scriptscriptstyle\pm 0.06}$
Ours $73.39_{\scriptscriptstyle\pm 0.12}$ $72.34_{\scriptscriptstyle\pm 0.13}$ $69.44_{\scriptscriptstyle\pm 0.17}$ $55.92_{\scriptscriptstyle\pm 0.02}$
ViT-B-16 Random $59.09_{\scriptscriptstyle\pm 0.49}$ $51.65_{\scriptscriptstyle\pm 0.46}$ $40.19_{\scriptscriptstyle\pm 0.13}$ $22.13_{\scriptscriptstyle\pm 0.16}$
Ours-LF $57.50_{\scriptscriptstyle\pm 0.69}$ $49.81_{\scriptscriptstyle\pm 0.89}$ $40.33_{\scriptscriptstyle\pm 0.86}$ $23.08_{\scriptscriptstyle\pm 0.27}$
Ours $57.62_{\scriptscriptstyle\pm 0.56}$ $52.67_{\scriptscriptstyle\pm 0.49}$ $42.15_{\scriptscriptstyle\pm 0.81}$ $24.43_{\scriptscriptstyle\pm 0.28}$
C.5 Algorithm for stratified sampling using CCS [86]
Here we present the algorithm for sampling the training examples to form the coreset based on the coverage-based selection methodology proposed by [86]. A crucial component of the algorithm is the cutoff rate $\beta$ which controls how many challenging samples should be removed from consideration when selecting the coreset. This is done to eliminate misclassified samples from the dataset since they can hurt the performance of the model trained on coreset, especially at high pruning rates. Previous works [86, 87] ablate the values of this cutoff ratio by training the downstream model on a range of values. In our work, we simply use the values proposed by the previous works and find that they work well for our score as well. The cutoff rates $\beta$ for different pruning rates $\alpha$ are as follows ($\alpha$, $\beta$). For CIFAR-10: (30%, 0), (50%, 0), (70%, 10%), (90%, 30%), for CIFAR-100: (30%, 10%), (50%, 20%), (70%, 20%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%). We used CCS for label-free CS as well and the cutoff rates used were for CIFAR-10: (30%, 0), (50%, 0), (70%, 20%), (90%, 40%), for CIFAR-100: (30%, 0), (50%, 20%), (70%, 40%), (90%, 50%), for Imagenet: (30%, 0), (50%, 10%), (70%, 20%), (90%, 30%).
Algorithm 1 Coverage-centric Coreset Selection (CCS) [86]
Input: Dataset with difficulty scores: $\mathbb{D}=\{(x,y,s)\}_{i=1}^{n}$, pruning ratio: $\alpha$, cutoff rate: $\beta$, number of bins: $b$.
Output: Coreset: $\mathcal{S}$
  # Prune hardest examples
[…]
  $A_{1},A_{2},\cdots,A_{b}\leftarrow$ Split scores in $\mathbb{D}^{\prime}$ into $b$ bins.
  $\mathcal{B}\leftarrow\{B_{i}:B_{i}\;\mathrm{consists\;of\;samples\;with\;scores\;in\;}A_{i}\;for\;i=1,\cdots,b\}$.
  # Define the size of the coreset
  $m\leftarrow n\times\alpha$.
  while $\mathcal{B}\neq\varnothing$ do
     # Select the bin with the fewest examples
     $B_{min}\leftarrow\arg\min_{B\in\mathcal{B}}|B|$.
     # Compute the budgets for this bin
     $m_{B}\leftarrow\min\{|B|,\left\lfloor\frac{m}{|\mathcal{B}|}\right\rfloor\}$.
     $\mathcal{S}_{B}\leftarrow$ randomly sample $m_{B}$ samples from $B_{min}$.
     $\mathcal{C}\leftarrow\mathcal{C}\bigcup\mathcal{S}_{B}.$
[…]
     $m\leftarrow m-m_{B}.$
  end while
  return $\mathcal{C}$.
C.6 Experimental details
For generating the importance score we pre-compute the concept similarity scores for the entire dataset and then train the concept-bottleneck layer (in block 2 of Fig. 2) for $100$ epochs across all experiments. This training only requires 800 seconds for Imagenet which is significantly more efficient than training the ResNet-34 model on Imagenet (requires roughly 8 hours on two A-100 GPUs). The accuracies of the models trained on the entire training set are $95.44$% and $78.74$% for ResNet(RN)-18 on CIFAR-10/100 and $72.4$% for RN-18, $75$% for RN-34, and $78.4$% for RN-50 on Imagenet.
After the coresets are selected, we use the setting and code from [86] for training a ResNet-18 model for $40000$ iterations with a batch size of $256$ on the coresets for all pruning ratios for CIFAR-10/CIFAR-100. For Imagenet, we train ResNet-18, ResNet-34, and ResNet-50 models for $100$ epochs on the coresets identified by our method using the training code based on FFCV [37].
The performance of the label-free CS is dependent on the quality of the pseudo-labels. Compared to the clustering-based approach used by ELFS [87], our approach of using the zero-shot classification ability of CLIP models yields significantly better pseudo-label quality along with being simpler and more efficient to compute. Specifically, for CIFAR-10/100, pseudo-labels of the training set are computed using the CLIP L-14 model trained on the DataComp-1B dataset [27] yields an accuracy of $98.52$% and $87.28$% whereas for Imagenet it achieves an accuracy of $79.47$% which are better than the best pseudo-label accuracy obtained by the clustering approach in ELFS ($92.5$% and $66.3$% on CIFAR-10/100 and $58.8$% on Imagenet).
For training the concept bottleneck layer we minimized the cross entropy loss using SGD with a learning rate of 1E-3, momentum of $0.9$ and a weight decay of 5E-4 for $100$ epochs.
(a) Easy images from CIFAR-10
(b) Challenging images from CIFAR-10
(c) Easy images from CIFAR-100
(d) Challenging images from CIFAR-100
Figure 4: Class-wise easy and challenging images for the 10 classes (airplane, car, bird, cat, deer, dog, frog, horse, ship, truck) in CIFAR-10 and for a subset of 10 classes (boy, bridge, camel, cloud, crab, kangaroo, lamp, rose, tiger, train) from CIFAR-100. Similar to the results in Fig. 3, easy images (a,c) are more canonical images associated with the class labels whereas challenging images (b,d) are images that are confused between two or more classes in the dataset.